### v2

# Distributed Solving of Linear Quadratic Optimal Controller with Terminal State Constraint

Abstract
This paper is concerned with the linear quadratic (LQ) optimal control of continuous-time system with terminal state constraint. In particular, multiple agents exist in the system which can only access partial information of the matrix parameters. This makes the classical solving method based on Riccati equation with global information suffering. The main contribution is to present a distributed algorithm to derive the optimal controller which is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state. The effectiveness of the proposed algorithm is verified by two numerical examples.
Index Terms:
LQ optimal control, Terminal state constraint, Distributed algorithm.
I Introduction
Linear quadratic (LQ) optimal control is one of the core fundamental problem in modern control theory which has wide applications in plenty of fields including power system scheduling [1], economics [2] and autonomous driving [3]. A typical example is the trajectory tracking control problem which can be solved by designing a LQ-based tracking controller [4]. For the classic LQ optimal control problem, one of the most powerful tools is Riccati equation [5], which can be used to design the optimal feedback gain. For the irregular LQ optimal control, a hierarchical optimization method has been proposed in [6] where two Riccati equations are given to design the optimal controller.
Considering the wide existence of terminal state constraints in lots of practical problems such as the precise terminal guidance in the hypersonic vehicle re-entry tracking control [7], it is necessary to investigate the LQ optimal control problems with terminal state constraints. [8] studied the nonlinear systems with terminal constraints by using a robust receding horizon controller. [9] considered a time-varying set of terminal constraints by using robust constrained model predictive control algorithm. [10] studied LQ optimal control with fixed terminal states and integral quadratic constraints. For the cases with stochastic uncertainty, [11] studied a general class of stochastic LQ control problems with terminal states constrained to convex sets. [12] investigated a stochastic LQ optimal control with a fixed terminal state and linear manifold constraint on the initial state. [13] obtained the explicit optimal solution for discrete-time stochastic LQ optimal control with initial and terminal state constraints.
Note that the design of the optimal controllers in the above mentioned results relies on the global information of the matrix parameters in the systems. This loses efficacy when there exist sensor networks composed of a large number of sensor and actuator nodes/agents with communication capabilities among them [14], which have received a great deal of attention in the past decade due mainly to its current and future envisioned applications including mobile vehicles, smart grids, environmental monitoring and industrial process control [15, 16, 17, 18]. In this case, each agent can only access partial information of the system and make decisions by using their own and/or neighbors’ information, which has given rise to decentralized and distributed control problems.
Decentralized control refers to situations where different control posts access different information and make decisions by using own information. [19] provided the optimal solutions to decentralized linear-quadratic-Gaussian (LQG) control problems with one-step delayed information sharing pattern by using dynamic programming. [20] derived the optimal linear solution of the decentralized LQG control with $d$-step delayed information sharing pattern. The partially nested decentralized LQG problems with controllers connected over a graph were studied in [21]. [22] studied teams with only control sharing, [23] studied the broadcast information structure, and [24] investigated the systems with common and private observations.
Distributed control refers to situations where different agents access different information, share their information through network topology, and make decisions by using own and neighbors’ information. One of the most popular problems is consensus control which has been well studied [25, 26]. For the optimal control problems, [27] studied the distributed optimal control for discrete-time systems. [28] discussed the distributed consensus optimal control strategy for multi-agent systems with a general directed graph. [29] investigated a globally optimal distributed formation control algorithm. [30] considered the online distributed convex optimization problem with local dynamic loss function sums under unbalanced graphs. A distributed LQG optimal control problem was investigated in [31], where each controller has its own input matrix.
Although significant progress has been made for the decentralized and distributed optimal control problem, the following issues still exist:
  • The matrix parameters in the systems are assumed to be completely known. While the case that each agent can only access partial information of the matrix parameters remains to be solved.
  • Strong conditions are imposed on the topology and/or the weighting matrices in the cost function to guarantee the distributed structure of the optimal solution, apart from the connectivity. While when these conditions are not established, how to derive the optimal solution in a distributed way remain to be studied.
  • No terminal state has been imposed in the distributed optimal control problem. While the existence of the terminal constraint additionally introduces an optimal Lagrange multiplier and a backward differential equation driven by the multiplier, whose distributed solving needs further study.
In this paper, we will study the LQ optimal control problem with terminal state constraint where there exist multiple agents and each agent accesses partial information of the matrix parameters. The main contribution is to present a distributed algorithm to derive the optimal controller. Specifically, the distributed algorithm is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state.
The remainder of the paper is organized as follows. Section II describes the studied problem. The explicit and the numerical solutions in centralized way are provided in Section III. Section IV gives the distributed algorithm for the optimal controller. Numerical example is given in Section V. The conclusion is outlined in the last Section. Detailed proofs of the results are given in Appendix.
II Problem formulation
Consider the following continuous-time system:
$\dot{x}(t)=Ax(t)+Bu(t),$ (1)
where $x\in R^{n}$ is the state, $u\in R^{m}$ is the control input. $A$ and $B$ are constant matrices with compatible dimensions. $t\in[0,T]$ is the continuous-time variable. The initial state is prescribed as $x(0)=x_{0}$. The terminal state is predetermined as $x(T)=x_{T}$.
The cost function is given by
[…]
where $Q$ is a positive semi-definite matrix with compatible dimension, and $R$ is a positive-definite matrix with compatible dimension.
In this paper, we study a novel case with a sensor network in the system that consists of $N$ agents or nodes, and the communication network between the agents is represented by the undirected topological graph $\mathscr{G}=\{\mathscr{V},\mathscr{E},\mathscr{A}\}$. In particular, each agent $i,i=1,2,\ldots,N$ can only communicate with its neighbors $N_{i}$ without needing global topology information. The graph is assumed to be connected. For more details about the graph refer to [28].
In this case, the matrix parameters $A,B$ can not be all available to the controller design of agent $i$. That is, different partial information is accessible to different agent which makes the design of the optimal controller much more complicated. In details, the following assumptions are given to show the detailed information available to agent $i$.
Assumption 1
There exist matrices $A_{i},B_{i},i=1,\ldots,N$ such that $A=\frac{1}{N}\sum_{i=1}^{N}A_{i}$, $B=\frac{1}{N}\sum_{i=1}^{N}B_{i}$. In particular, $A_{i}$ and $B_{i}$ are accessible for agent $i,i=1,\ldots,N$.
Assumption 2
There exist matrices $Q_{i},i=1,\ldots,N$ such that $Q=\frac{1}{N}\sum_{i=1}^{N}Q_{i}$, and $Q_{i}$ is accessible for agent $i,i=1,\ldots,N$. The matrix $R$ is accessible to each agent.
Assumption 3
There exist $x_{i,0},x_{i,T},i=1,\ldots,N$ such that $x(0)=x_{0}=\frac{1}{N}\sum_{i=1}^{N}x_{i,0}$ and $x(T)=x_{T}=\frac{1}{N}\sum_{i=1}^{N}x_{i,T}$. In particular, $x_{i,0},x_{i,T}$ accessible to agent $i,i=1,\ldots,N$.
Assumption 4
There exists matrices $M_{i},i=1,\ldots,N$ such that $BR^{-1}B^{\prime}=\frac{1}{N}\sum_{i=1}^{N}M_{i}^{\prime}M_{i}$, where $M_{i}$ is accessible for agent $i,i=1,\ldots,N$.
Based on the above description, the studied problem is concluded as follows.
Problem. The aim of the paper is to design a distributed algorithm for the optimal controller which minimizes the cost function (2) subject to system (1) and ensures the state to satisfy $x(0)=x_{0}$ and $x(T)=x_{T}$ for each agent $i$ under Assumptions 1-4.
Figure 1. The structural diagram of the Problem.
Remark 1
The structural diagram of the studied Problem is shown in Figure 1, where agent $i,i=1,2,\ldots,N$ can only access its own $A_{i}$, $B_{i}$ and communicate with its neighbors. The problem has important applications in the fields of urban intelligent surveillance, disaster emergency response, and collaborative search and rescue. For example, in the distributed control problem of heterogeneous multi-agent systems (HMASs) consisting of unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs) [32], the dynamics of the $i$ agent can be described by
$\dot{x}_{i}(t)=\mathcal{A}_{i}x_{i}(t)+\mathcal{B}_{i}u_{i}(t),i=1,..,N.$ (3)
where $x_{i}(t)\in R^{n}$, $u_{i}(t)\in R^{m}$. By defining $x(t)=[x_{1}(t)^{\prime}~{}\cdots~{}x_{N}(t)^{\prime}]^{\prime}$, $u(t)=[u_{1}(t)^{\prime}~{}\cdots~{}u_{N}(t)^{\prime}]^{\prime}$, the system (3) can be expressed in the global form of system (1) with $A=\frac{1}{N}\sum_{i=1}^{N}A_{i}$, $B=\frac{1}{N}\sum_{i=1}^{N}B_{i}$, where
$A_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&\mathcal{A}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],B_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&\mathcal{B}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right].$ (12)
It is obvious that each agent makes a decision based on its local information $A_{i}$, $B_{i}$, which aligns with the problem studied in this paper.
III Preliminaries
III-A Explicit solution by using all the information
In this part, we present the optimal solution to the Problem by assuming all the information of matrix parameters are available for the agents. Noting that the terminal state is predetermined, thus it is necessary to ensure that the system can be driven from the initial state to the terminal state. To this end, we firstly present the definition of the reachability of system (1) as below.
Definition 1: A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
Next, we introduce the Riccati differential equation:
$\dot{P}(t)+A^{\prime}P(t)+P(t)A+Q-P(t)BR^{-1}B^{\prime}P(t)=0,$ (13)
with terminal value $P(T)=0$ and the following matrix differential equation:
$\dot{\Phi}(t)=$ $~{}[A-BR^{-1}B^{\prime}P(t)]\Phi(t),$ (14)
with initial value $\Phi(0)=I$. Furthermore, we denote $\Phi(t,s)=\Phi(t)\Phi^{-1}(s)$ and $\Psi(t)={\Phi^{\prime}(t)}^{-1}$.
The optimal solution is then given as follows.
Theorem 1
If the terminal state $x_{T}$ is reachable from the initial state $x_{0}$ of system (1), then there exists a solution $\lambda^{*}$ to the following equation:
[…]
Then, the optimal controller of minimizing (2) subject to (1) with $x(0)=x_{0}$ and $x(T)=x_{T}$ is given by:
$u^{*}(t)=R^{-1}B^{\prime}P(t)x^{*}(t)-R^{-1}B^{\prime}\beta(t),$ (16)
where $\beta(t)$ obeys the following equation:
$\dot{\beta}(t)=-[A^{\prime}-P(t)BR^{-1}B^{\prime}]\beta(t),~{}\beta(T)=\lambda^{*}$ (17)
and $x^{*}(t)$ obeys the following equation:
$\dot{x}^{*}(t)$ $=$ $[A-BR^{-1}B^{\prime}P(t)]x^{*}(t)-BR^{-1}B^{\prime}\beta(t),$ (18)
with $x^{*}(0)=x_{0}$.
Proof: The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
III-B Numerical solution by using all the information
From Theorem 1, the design of optimal controller (16) depends on the solution $P(t)$ of the Riccati differential equation (13), which is nonlinear and typically challenging to solve analytically. To address this, we propose the following iteration:
$\dot{P}^{n}(t)$ $=-Z^{n}(t)^{\prime}P^{n}(t)-P^{n}(t)Z^{n}(t)-V^{n}(t),$ (19)
where $t\in[0,T]$ and
$Z^{n}(t)$ $=$ $A-BR^{-1}B^{\prime}P^{n-1}(t),$ (20)
$V^{n}(t)$ $=$ $Q+P^{n-1}(t)BR^{-1}B^{\prime}P^{n-1}(t),$ (21)
with $P^{n}(T)=0$, $P^{0}(t)=0$, and give its solution by
[…]
where $\Phi^{n}(t)$ and $\Psi^{n}(t)$ satisfy the following iteration respectively:
$\dot{\Phi}^{n}(t)$ $=$ $Z^{n}(t)\Phi^{n}(t),$ (23)
$\dot{\Psi}^{n}(t)$ $=$ $-Z^{n}(t)^{\prime}\Psi^{n}(t),$ (24)
with $\Phi^{n}(0)=I$ and $\Psi^{n}(0)=I$.
It is then easy to verify the convergence of the iteration (19) as given below.
Lemma 1
It holds that
(1) $P^{n}(t)\geq P^{n+1}(t)$;
(2)$\lim_{n\rightarrow\infty}P^{n}(t)=P(t)$, where $P(t)$ satisfies (13).
Proof: The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
Furthermore, it is immediate to obtain the following result.
Corollary 1
It holds that
$\lim_{n\rightarrow\infty}Z^{n}(t)$ $=$ $Z(t),$ (25)
$\lim_{n\rightarrow\infty}\Phi^{n}(t)$ $=$ $\Phi(t),$ (26)
$\lim_{n\rightarrow\infty}\Psi^{n}(t)$ $=$ $\Psi(t),$ (27)
where $Z(t)=A-BR^{-1}B^{\prime}P(t)$, $\Phi(t)$ satisfies (14), and $\Psi(t)=\Phi^{\prime}(t)^{-1}$.
Proof: The result follows directly from Lemma 1. $\blacksquare$
From Lemma 1 and Corollary 1, we have obtained the numerical solution to (13) and (14). Accordingly, the optimal solution (16) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
Algorithm 1 Centralized Algorithm
Input: Let the initial matrices be $P^{n}(T)=0,n=1,2,3,\ldots$, $P^{0}(t)=0,t\in[0,T]$, and select a sufficient small constant $\varepsilon>0$.
Output: The optimal states $x_{c}^{*}(t)$ and controller $u^{*}_{c}(t)$.
1. For $n=1,2,3,\ldots,$ run the iteration (19) and obtain $P^{n}(t)$.
2. If $\|P^{n}(t)-P^{n-1}(t)\|<\varepsilon$, end the iteration, and derive $P^{n}(t)$, $Z^{n}(t)$ and $\Phi^{n}(t)$.
3. Derive $\lambda_{c}^{*}$ by using
[…]
4. Solve the equation
$\dot{\beta}_{c}(t)=-Z^{n}(t)\beta_{c}(t),\beta_{c}(T)=\lambda_{c}^{*},$
and obtain $\beta_{c}(t)$.
5. Solve the equation
$\dot{x}_{c}^{*}(t)=Z^{n}(t)x^{*}(t)-BR^{-1}B^{\prime}\beta_{c}(t),x_{c}^{*}(0)=x_{0},$
and obtain $x_{c}^{*}(t)$.
6. Let $u_{c}^{*}(t)=-R^{-1}B^{\prime}P^{n}(t)x_{c}^{*}(t)-R^{-1}B^{\prime}\beta_{c}(t)$ and output $u_{c}^{*}(t)$ and $x_{c}^{*}(t)$.
Theorem 2
There exist an integer $n$ and a sufficient small constants $\epsilon_{1}>0$, $\epsilon_{2}>0$ such that output $x_{c}^{*}(t)$, $u_{c}^{*}(t)$ derived by Algorithm 1 satisfies
$\|x_{c}^{*}(t)-x^{*}(t)\|$$\leq$ $\epsilon_{1},$ (28)
$\|u_{c}^{*}(t)-u^{*}(t)\|$$\leq$ $\epsilon_{2},$ (29)
where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (16) and (18) by assuming all the information of system parameters are available to make decision.
Proof: From Lemma 1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$. This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
$\|\lambda_{c}-\lambda^{*}\|\leq\varepsilon_{1},$
$\|\beta_{c}(t)-\beta(t)\|\leq\varepsilon_{2},$
Accordingly, there exists a sufficient small constant $\epsilon_{1}>0$, $\epsilon_{2}>0$ such that $x_{c}^{*}(t)$ and $u_{c}^{*}(t)$ obtained by running Algorithm 1 satisfies (28) and (29). This completes the proof. $\blacksquare$
Remark 2
In Algorithm 1, solving the Riccati equation (19) requires $Z^{n}(t)$ and $V^{n}(t)$, which depends on the global system matrices $A$ and $B$. However, this loses efficacy if each agent only has access to the local information as defined in Assumptions 1-4. Something similar happens to $\lambda_{c}^{*}$, $x_{c}^{*}(t)$ and $\beta_{c}(t)$. This motivates the study of the distributed algorithm.
IV Main results
In this section, we will present a distributed algorithm to derive optimal controller when only partial information is accessible to agent $i$ as provided in Assumptions 1-4 which is divided into three steps. The first step is to give the distributed iteration for $P(t)$ in (13). The second step is to give the distributed iteration for optimal Lagrange multiplier $\lambda^{*}$ in (15) and the distributed iteration for $\beta(t)$ in (17). The third step is to give the distributed iteration for $x^{*}(t)$ in (18).
Step 1: Distributed iteration solution for $P(t)$
Firstly, we introduce the following distributed iteration of $P(t)$:
$\dot{P}_{i,k}^{n}(t)$ $=$ $-Z_{i,k}^{n}(t)^{\prime}P_{i,k}^{n}(t)-P_{i,k}^{n}(t)Z_{i,k}^{n}(t)-V_{i,k}^{n}(t),$ (30)
where $n=1,2,\cdots$, $k=1,2,\cdots$, $P_{i,0}^{n}(t)=0$, $P_{i.k}^{n}(T)=0$,
$Z_{i,k}^{n}(t)$ $=$ $Z_{i,k-1}^{n}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$ (31)
$-Z_{i,k-1}^{n}(t)]+\sum_{j\in N_{i}}[Z_{j,k-1}^{n}(t)-Z_{i,k-1}^{n}(t)],$
$V_{i,k}^{n}(t)$ $=$ $V_{i,k-1}^{n}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{n-1}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$ (32)
$-V_{i,k-1}^{n}(t)]+\sum_{j\in N_{i}}[V_{j,k-1}^{n}(t)-V_{i,k-1}^{n}(t)],$
while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
By solving (30), the explicit solution of $P_{i,k}^{n}(t)$ is
$P_{i,k}^{n}(t)$ $=$ $\Psi_{i,k}^{n}(t)\big{[}\int_{t}^{T}\Phi_{i,k}^{n}(s)^{\prime}V_{i,k}^{n}(s)$ (33)
×$\times\Phi_{i,k}^{n}(s)ds\big{]}\Psi_{i,k}^{n}(t)^{\prime}.$
where $\Phi_{i,k}^{n}(t)$ and $\Psi_{i,k}^{n}(t)$ satisfy the following iteration:
$\dot{\Phi}_{i,k}^{n}(t)$ $=$ $Z_{i,k}^{n}(t)\Phi_{i,k}^{n}(t),$ (34)
$\dot{\Psi}_{i,k}^{n}(t)$ $=$ $-Z_{i,k}^{n}(t)^{\prime}\Psi_{i,k}^{n}(t),$ (35)
with $\Phi_{i,k}^{n}(0)=I$ and $\Psi_{i,k}^{n}(0)=I$.
It is clear that the convergence of distributed iteration (30) depends on the convergence of distributed iteration (31)-(32) which is shown in the following result.
Lemma 2
It holds that
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|Z_{i,k}^{n}(t)-Z(t)\|$ $=$ $0,$ (36)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|V_{i,k}^{n}(t)-V(t)\|$ $=$ $0$ (37)
Proof: The detailed proof is given in Appendix A. $\blacksquare$
Then the following convergence result follows for the distributed iteration (30), (34) and (35).
Corollary 2
It holds that
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|P_{i,k}^{n}(t)-P(t)\|$ $=$ $0,$ (38)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Phi_{i,k}^{n}(t)-\Phi(t)\|$ $=$ $0,$ (39)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Psi_{i,k}^{n}(t)-\Psi(t)\|$ $=$ $0$ (40)
Proof: The result follows directly from Lemma 2. $\blacksquare$
For the convenience of marking, we make the denotations:
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}P_{i,k}^{n}(t)=P^{\infty}_{i}(t),$ (41)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\Phi_{i,k}^{n}(t)=\Phi^{\infty}_{i}(t),$ (42)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\Psi_{i,k}^{n}(t)=\Psi^{\infty}_{i}(t),$ (43)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}Z_{i,k}^{n}(t)=Z^{\infty}_{i}(t),$ (44)
$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}V_{i,k}^{n}(t)=V^{\infty}_{i}(t).$ (45)
Step 2: Distributed iteration solutions for $\lambda^{*}$ and $\beta(t)$
Secondly, to derive the distributed solution of $\lambda^{*}$, we denote
[…]
where $W=R^{-\frac{1}{2}}B^{\prime}$ and introduce the following distributed iteration for matrix $W$ as:
$W_{i,k}$ $=$ $W_{i,k-1}+\alpha_{k}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,k-1}]$ (47)
$+\sum_{j\in N_{i}}[W_{j,k-1}-W_{i,k-1}],$
whose convergence is given below.
Lemma 3
It holds that
$W_{i}^{\infty}\triangleq\displaystyle\lim_{k\rightarrow\infty}W_{i,k}=W.$ (48)
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
Recalling (15) where $\lambda^{*}$ depends not only on the global $W$ but also on the global initial state $x(0)$ and terminal state $x(T)$, we then propose the distributed iteration of $\lambda^{*}$ by
$\lambda_{i,k}^{*}$ $=$ $\lambda_{i,k-1}^{*}+\alpha_{k}\{(\varrho_{i}^{\infty})^{{\dagger}}[\Phi^{\infty}_{i}(T,0)x_{i,0}-x_{i,T}]$ (49)
$-\lambda_{i,k-1}^{*}\}+\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
where
[…]
and show the convergence in the following result.
Lemma 4
It holds that
$\lambda_{i}^{\infty,*}\triangleq\displaystyle\lim_{k\rightarrow\infty}\lambda_{i,k}^{*}=\lambda^{*}.$ (51)
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
Accordingly, the distributed solution for $\beta(t)$ is defined as follows:
$\dot{\beta}_{i}^{\infty}(t)$ $=$ $-Z^{\infty}_{i}(t)\beta_{i}^{\infty}(t),$ (52)
$\beta_{i}^{\infty}(T)$ $=$ $\lambda_{i}^{\infty,*}.$
Step 3: Distributed iteration solution for $x^{*}(t)$
Thirdly, we give the distributed solution of $x^{*}(t)$. In fact, by solving (15), the explicit solution of $x^{*}(t)$ is given by
[…]
which relies on the global initial state $x_{0}$. To avoid the using of global information, we give the distributed iteration as follows:
$x_{i,k}^{*}(t)$ $=$ $x_{i,k-1}^{*}(t)+\alpha_{k}[\Phi_{i}^{\infty}(t,0)x_{i,0}-\int_{0}^{t}\Phi_{i}^{\infty}(t,s)$ (54)
×$\times{W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\beta_{i}^{\infty}(s)ds-x_{i,k-1}^{*}(t)]$
$+\sum_{j\in N_{i}}[x_{j,k-1}^{*}(t)-x_{i,k-1}^{*}(t)],$
and derive the following convergence result.
Lemma 5
It holds that
$\displaystyle\lim_{k\rightarrow\infty}x_{i,k}^{*}(t)=x^{*}(t).$ (55)
Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
From Lemmas 2-5 and Corollary 2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
Algorithm 2 Distributed Algorithm
Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$, $P_{i,k}^{0}(t)=0$, $W_{i,0}=0$, $\lambda_{i,0}^{*}=0$, and $x_{i,0}^{*}(t)=0$ and select a sufficient small constant $\varsigma>0$.
Output: The optimal states $x_{i,w}^{*}(t)$ and distributed controller $u_{i}^{*}(t)$, $i=1,2,\ldots,N$.
  1. 1.
    For $n=1,2,\cdots$,
    • For $k=1,2,3,\cdots$, run the iteration (31)-(32) and obtain $Z_{i,k}^{n}(t)$, $V_{i,k}^{n}(t)$. If $\|Z_{i,k}^{n}(t)-Z_{i,k-1}^{n}(t)\|<\varsigma$ and $\|V_{i,k}^{n}(t)-V_{i,k-1}^{n}(t)\|<\varsigma$, end the iteration, and derive $Z_{i,k}^{n}(t)$ and $V_{i,k}^{n}(t)$.
    • Solve the equation (30) and obtain $P_{i,k}^{n}(t)$. If $\|P_{i,k}^{n}(t)-P_{i,k}^{n-1}(t)\|<\varsigma$, end the iteration, and derive $P^{n}_{i,k}(t)$, $\Phi^{n}_{i,k}(t)$, $\Psi^{n}_{i,k}(t)$.
  2. 2.
    For $\varpi=1,2,\cdots$, run the iteration
    $W_{i,\varpi}$ $=$ $W_{i,\varpi-1}+\alpha_{\varpi}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,\varpi-1}]$
    $+\sum_{j\in N_{i}}[W_{j,\varpi-1}-W_{i,\varpi-1}].$
    • If $\|W_{i,\varpi}-W_{i,\varpi-1}\|<\varsigma$, end the iteration, and derive $W_{i,\varpi}$.
  3. 3.
    For $q=1,2,\cdots$, run the iteration
    $\lambda_{i,q}^{*}$ $=$ $\lambda_{i,q-1}^{*}+\alpha_{q}\bigg{\{}\Big{[}\int_{0}^{T}\Phi_{i,k}^{n}(s)W_{i,\varpi}^{\prime}W_{i,\varpi}$
    ×$\times\Psi_{i,k}^{n}(s)ds\Big{]}^{{\dagger}}[\Phi_{i,k}^{n}(T,0)x_{i,0}-x_{i,T}]$
    $-\lambda_{i,q-1}^{*}\}+\sum_{j\in N_{i}}[\lambda_{j,q-1}^{*}-\lambda_{i,q-1}^{*}].$
    • If $\|\lambda_{i,q}^{*}-\lambda_{i,q-1}^{*}\|<\varsigma$, end the iteration, and derive $\lambda_{i,q}^{*}$.
  4. 4.
    Solve the equation
    $\dot{\beta}_{i}(t)=-Z_{i,k}^{n}(t)\beta_{i}(t),~{}\beta_{i}(T)=\lambda_{i,q}^{*},$
    and obtain $\beta_{i}(t)$.
  5. 5.
    For $w=1,2,\cdots$, run the iteration
    $x_{i,w}^{*}(t)$ $=$ $x_{i,w-1}^{*}(t)+\alpha_{w}[\Phi_{i,k}^{n}(t,0)x_{i,0}-\int_{0}^{t}\Phi_{i,k}^{n}(t,s)$
    ×$\times W_{i,\varpi}^{\prime}W_{i,\varpi}\beta_{i}(s)ds-x_{i,w-1}^{*}(t)]$
    $+\sum_{j\in N_{i}}[x_{j,w-1}^{*}(t)-x_{i,w-1}^{*}(t)].$
    • If $\|x_{i,w}^{*}(t)-x_{i,w-1}^{*}(t)\|<\varsigma$, end the iteration, and derive $x_{i,w}^{*}(t)$.
  6. 6.
    Let $u_{i}^{*}(t)=-R^{-1/2}W_{i,\varpi}^{\prime}\big{[}P_{i,k}^{n}(t)x_{i,w}^{*}(t)+\beta_{i}(t)\big{]}$ and output $u_{i}^{*}(t)$ and $x_{i,w}^{*}(t)$.
Theorem 3
There exist integers $n,k,\varpi,q,w$ and sufficiently small constants $\psi_{1}>0$, $\psi_{2}>0$ such that output $x_{i,w}^{*}$ and $u_{i}^{*}(t)$ derived by Algorithm 2 satisfy:
$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (16) and (18) by assuming all the information of system parameters are available to make decision.
Proof. From Lemmas 1-2 and Corollary 2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$. Together with Lemmas 3-4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
Accordingly, by using Lemma 5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57). This completes the proof. $\blacksquare$
In the following, we generalize the above results to the standard LQ problem without terminal state constraint, i.e. minimizing (2) s. t. (1) with $x(0)=x_{0}$.
In this case, the optimal controller $u^{\diamond}(t)$ is
$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
where $P(t)$ is the solution of equation (13), $x^{\diamond}(t)$ is the solution of the following equation:
$\dot{x}^{\diamond}(t)$ $=$ $[A-BR^{-1}B^{\prime}P(t)]x^{\diamond}(t),~{}x^{\diamond}(0)=x_{0}.$ (59)
Based on Algorithm 2, the distributed design of the optimal controller (58) is shown in Algorithm 3, the convergence of which is given below.
Algorithm 3 Distributed Algorithm without Terminal State Constraint
Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{\bar{n}}(t)=0$, $P_{i,\bar{n}}^{0}(t)=0$, $W_{i,0}=0$, and $x_{i,0}^{\diamond}(t)=0$ and select a sufficient small constant $\bar{\varsigma}>0$.
Output: The distributed controller $u_{i}^{\diamond}(t)$, $i=1,2,\ldots,N$.
  1. 1.
    For $\bar{n}=1,2,\cdots$,
    • For $\bar{k}=1,2,3,\cdots$, run the iteration (31)-(32) and obtain $Z_{i,\bar{k}}^{\bar{n}}(t)$, $V_{i,\bar{k}}^{\bar{n}}(t)$. If $\|Z_{i,\bar{k}}^{n}(t)-Z_{i,\bar{k}-1}^{\bar{n}}(t)\|<\bar{\varsigma}$ and $\|V_{i,\bar{k}}^{n}(t)-V_{i,\bar{k}-1}^{\bar{n}}(t)\|<\bar{\varsigma}$, end the iteration, and derive $Z_{i,\bar{k}}^{\bar{n}}(t)$ and $V_{i,\bar{k}}^{\bar{n}}(t)$.
    • Solve the equation (30) and obtain $P_{i,\bar{k}}^{\bar{n}}(t)$. If $\|P_{i,\bar{k}}^{\bar{n}}(t)-P_{i,\bar{k}}^{\bar{n}-1}(t)\|<\bar{\varsigma}$, end the iteration, and derive $P^{\bar{n}}_{i,\bar{k}}(t)$, $\Phi^{\bar{n}}_{i,\bar{k}}(t)$.
  2. 2.
    For $\bar{\varpi}=1,2,\cdots$, run the iteration
    $W_{i,\bar{\varpi}}$ $=$ $W_{i,\bar{\varpi}-1}+\alpha_{\bar{\varpi}}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,\bar{\varpi}-1}]$
    $+\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
    • If $\|W_{i,\bar{\varpi}}-W_{i,\bar{\varpi}-1}\|<\bar{\varsigma}$, end the iteration, and derive $W_{i,\bar{\varpi}}$.
  3. 3.
    For $\bar{w}=1,2,\cdots$, run the iteration
    $x_{i,\bar{w}}^{\diamond}(t)$ $=$ $x_{i,\bar{w}-1}^{\diamond}(t)+\alpha_{\bar{w}}[\Phi_{i,\bar{k}}^{n}(t,0)x_{i}(0)-x_{i,\bar{w}-1}^{\diamond}(t)]$
    $+\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
    • If $\|x_{i,\bar{w}}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)\|<\bar{\varsigma}$, end the iteration, and derive $x_{i,\bar{w}}^{\diamond}(t)$.
  4. 4.
    Let $u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(t)x_{i,\bar{w}}^{\diamond}(t)$ and output $u_{i}^{\diamond}(t)$ and $x_{i,\bar{w}}^{\diamond}(t)$.
Corollary 3
There exist integers $\bar{n},\bar{k},\bar{\varpi},\bar{w}$ and sufficiently small constants $\bar{\psi}_{1}>0$, $\bar{\psi}_{2}>0$ such that output $x_{i,\bar{w}}^{\diamond}(t)$ and $u_{i}^{\diamond}(t)$ derived by Algorithm 3 satisfies
$\|x_{i,\bar{w}}^{\diamond}(t)-x^{\diamond}(t)\|$ $\leq\bar{\psi}_{1},$ (60)
$\|u_{i}^{\diamond}(t)-u^{\diamond}(t)\|$ $\leq\bar{\psi}_{2},$ (61)
where $x^{\diamond}(t)$ and $u^{\diamond}(t)$ are the optimal solution of (59) and (58) to the standard LQ problem without terminal state constraint by assuming all the information of system parameters are available to make decision.
Proof. The proof is similar to that of Theorem 3 and is thus omitted here.
Remark 3
The distributed solving of the finite-horizon solution (58) has been given in Algorithms 3. When the horizon $T$ is sufficiently large, the output of Algorithm 3
$u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(T)x_{i,\bar{w}}^{\diamond}(t).$ (62)
is an approximation solution to the standard infinite-horizon LQ problem under the assumptions of the stabilization of $(A,B)$ and observability of $(A,Q)$.
V Simulation Example
In this section, two examples are provided to show the effectiveness of the proposed Algorithms.
Example 1. Consider the distributed optimal control of a heterogeneous UGVs system, where the dynamics of each UGV is given by [35]
$\left[\begin{array}[]{c}\dot{q}_{i}(t)\\ \dot{v}_{i}(t)\end{array}\right]=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right]\left[\begin{array}[]{c}q_{i}(t)\\ v_{i}(t)\end{array}\right]+\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right]F_{i}(t),$ (63)
while $q_{i}(t)\in R^{2}$, $i=1,\cdots,5$, denote the position and velocity vectors of the $i$th UGV, $F_{i}(t)\in R^{2}$ denotes the force applied to the UGV, $D_{i}$ denotes the mechanical mass and $C_{i}$ denotes the translational friction coefficient. The communication topology among the UGVs is illustrated in Figure 2.
Figure 2. Communication topology among heterogeneous UGVs.
By defining the state variables as $x(t)=[x_{1}(t)^{\prime}~{}\cdots~{}x_{5}(t)^{\prime}]^{\prime}$ and the input variable as $u(t)=[u_{1}(t)^{\prime}~{}\cdots~{}u_{5}(t)^{\prime}]^{\prime}$, where ${x}_{i}(t)=[q_{i}(t)^{\prime}~{}~{}v_{i}(t)^{\prime}]^{\prime}$, ${u}_{i}(t)=F_{i}(t)$, then the UGV system (63) can be formulated as (3) with
$A_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&5\tilde{A}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],B_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&5\tilde{B}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],$
$\tilde{A}_{i}=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right],~{}~{}\tilde{B}_{i}=\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right].$
The control objective is to drive the UGVs from initial positions to predefined target formation while minimizing the cost functions (2) with $Q=16I$, $R=10I$, the initial states of the UGVs being
$x_{1}(0)=\left[\begin{array}[]{cccc}1.5\\ 2.5\\ 0\\ 0\end{array}\right],x_{2}(0)=\left[\begin{array}[]{cccc}2.5\\ 2.6\\ 0\\ 0\end{array}\right],x_{3}(0)=\left[\begin{array}[]{cccc}2\\ 1\\ 0\\ 0\end{array}\right],$
$x_{4}(0)=\left[\begin{array}[]{cccc}0\\ 1\\ 0\\ 0\end{array}\right],x_{5}(0)=\left[\begin{array}[]{cccc}1\\ 2\\ 0\\ 0\end{array}\right].$
and the terminal states of the UGVs being
$x_{1}(T)=\left[\begin{array}[]{cccc}5.5\\ 8\\ 3\\ 4\end{array}\right],x_{2}(T)=\left[\begin{array}[]{cc}7\\ 7\\ 2\\ 4\end{array}\right],x_{3}(T)=\left[\begin{array}[]{cc}8\\ 6\\ 3\\ 6\end{array}\right],$
$x_{4}(T)=\left[\begin{array}[]{cc}3\\ 6\\ 1\\ 3\end{array}\right],x_{5}(T)=\left[\begin{array}[]{cc}4\\ 7\\ 2\\ 1\end{array}\right].$
Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8. It is obvious that all the agents achieve consensus and converge to the centralized solutions from Figure 7-Figure 8. In particular, the trajectories of the agent achieve the predefined target formation at the given time $T=2$s as shown in Figure 9.
Figure 3. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
Figure 4. The trajectories of $\|W_{i,\varpi}\|_{2}$.
Figure 5. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
Figure 6. The trajectories of $\|\beta_{i}(t)\|_{2}$.
Figure 7. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$ with $w=200$.
Figure 8. The trajectories of $u_{i}^{*}(t)$.
Figure 9. The predefined target formation of heterogeneous UGVs.
Example 2. Consider system (1) and the cost function (2) with parameters given by
$A=\left[\begin{array}[]{cc}1{}&2\\ 0{}&1\end{array}\right],B=\left[\begin{array}[]{cc}1\\ 1\end{array}\right],Q=I,R=1,$ (74)
$x_{0}=\left[\begin{array}[]{cc}4\\ 4\end{array}\right],x_{T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right].$ (79)
Moreover, 4 agents are involved in the sensor network and the topology graph is shown in Figure 10.
Figure 10. Communication topology among agents.
Next, we consider the distributed case where each agent only has access to partial system information. Based on Assumptions 1-4, the parameters for each agent are given by
$A_{1}=\left[\begin{array}[]{cc}1.5{}&2\\ 0{}&0.5\end{array}\right],A_{2}=\left[\begin{array}[]{cc}0.5{}&2\\ 0{}&1.2\end{array}\right],A_{3}=\left[\begin{array}[]{cc}1.3{}&2\\ 0{}&1.8\end{array}\right],$
$A_{4}=\left[\begin{array}[]{cc}0.7{}&2\\ 0{}&0.5\end{array}\right],B_{1}=\left[\begin{array}[]{cc}1.5\\ 0.5\end{array}\right],B_{2}=\left[\begin{array}[]{cc}0.5\\ 1.2\end{array}\right],$
$B_{3}=\left[\begin{array}[]{cc}0.3\\ 1.8\end{array}\right],B_{4}=\left[\begin{array}[]{cc}1.7\\ 0.5\end{array}\right],Q_{1}=\left[\begin{array}[]{cc}1.2{}&0\\ 0{}&1\end{array}\right],$
$Q_{2}=\left[\begin{array}[]{cc}1{}&0\\ 0{}&0.8\end{array}\right],Q_{3}=\left[\begin{array}[]{cc}1{}&0\\ 0{}&1.2\end{array}\right],Q_{4}=\left[\begin{array}[]{cc}0.8{}&0\\ 0{}&1\end{array}\right],$
$R=1,x_{1,0}=\left[\begin{array}[]{cc}3.5\\ 4\end{array}\right],x_{2,0}=\left[\begin{array}[]{cc}4.5\\ 3.5\end{array}\right],x_{3,0}=\left[\begin{array}[]{cc}2.5\\ 4.5\end{array}\right],$
$x_{4,0}=\left[\begin{array}[]{cc}5.5\\ 4\end{array}\right],x_{i,T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right],i=1,\cdots,4.$
By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16. It is obvious that the distributed solutions converge to the centralized solutions Figure 15-Figure 16.
Figure 11. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
Figure 12. The trajectories of $\|W_{i,\varpi}\|_{2}$.
Figure 13. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
Figure 14. The trajectories of $\|\beta_{i}(t)\|_{2}$.
Figure 15. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$.
Figure 16. The trajectories of $u_{i}^{*}(t)$.
VI Conclusions
This paper studied the LQ optimal control problem for continuous-time systems with terminal state constraints. A distributed algorithm was proposed to solve the optimal controller where only partial information is accessible to the agent, consisting of the distributed iteration of the Riccati equation, the optimal Lagrange multiplier, and the optimal state. Numerical examples demonstrated the effectiveness of the proposed algorithm.
Appendix A
Proof of Lemma 2
We will prove the result by using mathematical induction. For $n=1$, it is obvious from (31)-(32) that
$Z_{i,k}^{1}(t)$ $=$ $Z_{i,k-1}^{1}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$ (86)
$-Z_{i,k-1}^{1}(t)]+\sum_{j\in N_{i}}[Z_{j,k}^{1}(t)-Z_{i,k-1}^{1}(t)],$
$V_{i,k}^{1}(t)$ $=$ $V_{i,k-1}^{1}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{0}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$ (87)
$-V_{i,k-1}^{1}(t)]+\sum_{j\in N_{i}}[V_{j,k}^{1}(t)-V_{i,k-1}^{1}(t)].$
It is aimed to show
$\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z(t)\|$ $=$ $0,$ (88)
$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V(t)\|$ $=$ $0$ (89)
which is divided into two steps. The first step is to prove (86)-(87) achieves consensus, that is,
$\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z_{j,k}^{1}(t)\|$ $=$ $0,$ (90)
$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V_{j,k}^{1}(t)\|$ $=$ $0$ (91)
The second step is to prove (86)-(87) achieves global convergence, that is,
$\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}Z_{i,k}^{1}(t)-Z^{1}(t)\|$ $=$ $0,$ (92)
$\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0$ (93)
Firstly, we prove that the consensus property by denoting
$\mathcal{Z}_{k}^{1}(t)$ $=$ $\textbf{col}\{Z_{1,k}^{1}(t),Z_{2,k}^{1}(t),\cdots,Z_{N,k}^{1}(t)\},$
$\mathcal{V}_{k}^{1}(t)$ $=$ $\textbf{col}\{V_{1,k}^{1}(t),V_{2,k}^{1}(t),\cdots,V_{N,k}^{1}(t)\},$
$\Pi_{k}^{1}(t)$ $=$ $\textbf{col}\{Y_{1,k}^{1}(t),Y_{2,k}^{1}(t),\cdots,Y_{N,k}^{1}(t))\},$
$\Theta_{k}^{1}(t)$ $=$ $\textbf{col}\{L_{1,k}^{1}(t),L_{2,k}^{1}(t),\cdots,L_{N,k}^{1}(t)\},$
$Y_{i,k}^{1}(t)$ $=$ $A_{i}-M_{i}M_{i}^{\prime}P_{i,k}^{0}(t)-Z_{i,k}^{1}(t),$
$L_{i,k}^{1}(t)$ $=$ $Q_{i}+P_{i,k}^{0}(t)M_{i}M_{i}^{\prime}P_{i,k}^{0}(t)-V_{i,k}^{1}(t),$
it follows from (86)-(87) that
$\mathcal{Z}_{k}^{1}(t)$ $=$ $(\mathcal{A}\otimes I_{n})\mathcal{Z}_{k-1}^{1}(t)+\alpha_{k}\Pi_{k-1}^{1}(t),$
$\mathcal{V}_{k}^{1}(t)$ $=$ $(\mathcal{A}\otimes I_{n})\mathcal{V}_{k-1}^{1}(t)+\alpha_{k}\Theta_{k-1}^{1}(t),$
where $\mathcal{A}=I_{N}-\mathcal{L}$, $\mathcal{L}$ is the Laplacian matrix [31]. It is obvious that $\mathcal{Z}_{k}^{1}(t)$ and $\mathcal{V}_{k}^{1}(t)$ are bounded. In fact, by defining $\bar{\mathcal{Z}}_{k}^{1}(t)=(U^{\prime}\otimes I_{n})\mathcal{Z}_{k}^{1}(t)$ and $\bar{\mathcal{V}}_{k}^{1}(t)=(U^{\prime}\otimes I_{n})\mathcal{V}_{k}^{1}(t)$, where $U$ is orthogonal matrix such that $U^{\prime}\mathcal{L}U=diag\{\lambda_{1},\cdots,\lambda_{N}\}$ and $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $\mathcal{L}$. According to (86)-(87), we have that
$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
$-\alpha_{k}(U^{\prime}\otimes I_{n})\Pi_{k-1}^{1}(t),$
$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
$-\alpha_{k}(U^{\prime}\otimes I_{n})\Theta_{k-1}^{1}(t).$
Together with the properties of Laplacian matrices [31], it is easy to get $\bar{\mathcal{Z}}_{k}^{1}(t)$ and $\bar{V}_{k}^{1}(t)$ are bounded. This implies the boundedness of $\mathcal{Z}_{k}^{1}(t)$ and $\mathcal{V}_{k}^{1}(t)$.
Furthemore, let $\mathcal{B}=\frac{1}{N}1_{N}1_{N}^{\prime}$, $\delta_{k}^{1}(t)=[(I_{N}-\mathcal{B})\otimes I_{n}]\mathcal{Z}^{1}_{k}(t)$, and $\eta_{k}^{1}(t)=[(I_{N}-\mathcal{B})\otimes I_{n}]\mathcal{V}_{k}^{1}(t)$, we have that
$\delta_{k}^{1}(t)$ $=$ $[(\mathcal{A}-\mathcal{B})\otimes I_{n}]\delta_{k-1}^{1}(t)$ (94)
$+\alpha_{k}[(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{k-1}^{1}(t),$
$\eta_{k}^{1}(t)$ $=$ $[(\mathcal{A}-\mathcal{B})\otimes I_{n}]\eta_{k-1}^{1}(t)$ (95)
$+\alpha_{k}[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{k-1}^{1}(t),$
where the fact $\mathcal{A}\mathcal{B}=\mathcal{B}\mathcal{A}=\mathcal{B}^{2}=\mathcal{B}$ has been used in the derivations. By applying iteration calculation to (94) and (95), it yields that
$\delta_{k}^{1}(t)$ $=$ $[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\delta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}$
×$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{\tau-1}^{1}(t),$
$\eta_{k}^{1}(t)$ $=$ $[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\eta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}$
×$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{\tau-1}^{1}(t).$
This further gives that
$\big{\|}\delta_{k}^{1}(t)\big{\|}$ $=$ $\big{\|}[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\delta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})$
$\otimes I_{n}]^{k-\tau}(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{\tau-1}^{1}(t)\big{\|}$
$\leq$ $c\rho^{k}\|\delta_{0}^{1}(t)\|+\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|,$
$\|\eta_{k}^{1}(t)\|$ $=$ $\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\eta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})$
$\otimes I_{n}]^{k-\tau}[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{\tau-1}^{1}(t)\|$
$\leq$ $c\rho^{k}\|\eta_{0}^{1}(t)\|+\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|.$
In view of the fact that the graph is connected, it yields that
$\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\leq c\rho^{k},$
where $c>0$ and $\rho\in(0,1)$. According to $\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|<\infty$, $\|\delta_{0}^{1}(t)\|<\infty$, $\|\eta_{0}^{1}(t)\|<\infty$, $\|\Pi_{\tau}^{1}(t)\|<\infty$, $\|\Theta_{\tau}^{1}(t)\|<\infty$, we have that
$\lim_{k\to\infty}c\rho^{k}\|\delta_{0}^{1}\|=0,$ (96)
$\lim_{k\to\infty}c\rho^{k}\|\eta_{0}^{1}\|=0,$ (97)
$\lim_{k\to\infty}\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
×$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|=0,$ (98)
$\lim_{k\to\infty}\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
×$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|=0.$ (99)
Thus, we obtain that
$\lim_{k\to\infty}\|\delta_{k}^{1}(t)\|$ $=$ $0,$
$\lim_{k\to\infty}\|\eta_{k}^{1}(t)\|$ $=$ $0,$
that is, (90)-(91) hold.
Secondly, we prove that (92)-(93) hold. To do this, we define $\tilde{Z}_{k}^{1}(t)=\frac{1}{N}\sum_{i=1}^{N}Z_{i,k}^{1}(t)$ and $\tilde{V}_{k}^{1}(t)=\frac{1}{N}\sum_{i=1}^{N}V_{i,k}^{1}(t)$, and derive from (86)-(87) that
$\tilde{Z}_{k}^{1}(t)$ $=$ $\tilde{Z}_{k-1}^{1}(t)+\frac{1}{N}\alpha_{k}\sum_{i=1}^{N}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$
$-Z_{i,k-1}^{1}(t)],$
$\tilde{V}_{k}^{1}(t)$ $=$ $\tilde{V}_{k-1}^{1}(t)+\frac{1}{N}\alpha_{k}\sum_{i=1}^{N}[Q_{i}+P_{i,k-1}^{0}(t)M_{i}M_{i}^{\prime}$
×$\times P_{i,k-1}^{0}(t)-V_{i,k-1}^{1}(t)].$
By defining $Z_{k}^{1}(t)$, $V_{k}^{1}(t)$ by
$Z_{k}^{n}(t)$ $=$ $Z_{k-1}^{n}(t)+\alpha_{k}[A-BR^{-1}B^{\prime}P^{n-1}(t)$ (100)
$-Z_{k-1}^{n}(t)],$
$V_{k}^{n}(t)$ $=$ $V_{k-1}^{n}(t)+\alpha_{k}[Q+P^{n-1}(t)BR^{-1}B^{\prime}$ (101)
×$\times P^{n-1}(t)-V_{k-1}^{n}(t)],$
and letting $\Delta Z_{k}^{1}(t)=\tilde{Z}_{k}^{1}(t)-Z_{k}^{1}(t)$, $\Delta V_{k}^{1}(t)=\tilde{V}_{k}^{1}(t)-V_{k}^{1}(t)$, we have that
$\Delta{Z}_{k}^{1}(t)$ $=$ $(1-\alpha_{k})\Delta{Z}_{k-1}^{1}(t),$
$\Delta{V}_{k}^{1}(t)$ $=$ $(1-\alpha_{k})\Delta{V}_{k-1}^{1}(t).$
By applying the facts that $10<1-\alpha_{k}<$ 1, and $\lim_{k\to\infty}\alpha_{k}=0$, it yields that
$\lim_{k\to\infty}\|\Delta{Z}_{k}^{1}(t)\|$ $=$ $0,$ (102)
$\lim_{k\to\infty}\|\Delta{V}_{k}^{1}(t)\|$ $=$ $0$ (103)
Combining with (90)-(91), it follows that
$\lim_{k\to\infty}\|Z_{i,k}^{1}(t)-Z_{k}^{1}(t)\|$ $=$ $0,$
$\lim_{k\to\infty}\|V_{i,k}^{1}(t)-V_{k}^{1}(t)\|$ $=$ $0$
Moreover, together with the fact that $\lim_{k\rightarrow\infty}\|Z_{k}^{1}(t)-Z^{1}(t)\|=0$ and $\lim_{k\rightarrow\infty}\|V_{k}^{1}(t)-V^{1}(t)\|=0$ obtained by using Theorem 3.1.1 in [36], we obtain (88)-(89) directly.
To complete the induction proof, assume that for any $m\leq n$, (36)-(37) holds, and $P_{i,k}^{m}(t)$ obtained by (30) is convergent. By taking similar proof for (90)-(91) and (92)-(93), it can be obtained that (36)-(37) holds when $n=m+1$. To avoid repetition, we omit the details of the proof here. $\blacksquare$

### v3

@@
 The effectiveness of the proposed algorithm is verified by two numerical examples.
-Index Terms:
+keywords:
 LQ optimal control, Terminal state constraint, Distributed algorithm.
-I Introduction
+thanks: Corresponding author
+1 Introduction
 Linear quadratic (LQ) optimal control is one of the core fundamental problem in modern control theory which has wide applications in plenty of fields including power system scheduling [1], economics [2] and autonomous driving [3].
@@
 Detailed proofs of the results are given in Appendix.
-II Problem formulation
+2 Problem formulation
 Consider the following continuous-time system:
@@
 In this paper, we study a novel case with a sensor network in the system that consists of $N$ agents or nodes, and the communication network between the agents is represented by the undirected topological graph $\mathscr{G}=\{\mathscr{V},\mathscr{E},\mathscr{A}\}$.
-In particular, each agent $i,i=1,2,\ldots,N$ can only communicate with its neighbors $N_{i}$ without needing global topology information.
+In particular, each agent $i,i=1,2,\ldots,N$ can only communicate with its neighbors $N_{i}$.
 The graph is assumed to be connected.
@@
 In details, the following assumptions are given to show the detailed information available to agent $i$.
-Assumption 1
+Assumption 1.
 There exist matrices $A_{i},B_{i},i=1,\ldots,N$ such that $A=\frac{1}{N}\sum_{i=1}^{N}A_{i}$, $B=\frac{1}{N}\sum_{i=1}^{N}B_{i}$.
 In particular, $A_{i}$ and $B_{i}$ are accessible for agent $i,i=1,\ldots,N$.
-Assumption 2
+Assumption 2.
 There exist matrices $Q_{i},i=1,\ldots,N$ such that $Q=\frac{1}{N}\sum_{i=1}^{N}Q_{i}$, and $Q_{i}$ is accessible for agent $i,i=1,\ldots,N$. The matrix $R$ is accessible to each agent.
-Assumption 3
+Assumption 3.
 There exist $x_{i,0},x_{i,T},i=1,\ldots,N$ such that $x(0)=x_{0}=\frac{1}{N}\sum_{i=1}^{N}x_{i,0}$ and $x(T)=x_{T}=\frac{1}{N}\sum_{i=1}^{N}x_{i,T}$.
 In particular, $x_{i,0},x_{i,T}$ accessible to agent $i,i=1,\ldots,N$.
-Assumption 4
+Assumption 4.
 There exists matrices $M_{i},i=1,\ldots,N$ such that $BR^{-1}B^{\prime}=\frac{1}{N}\sum_{i=1}^{N}M_{i}^{\prime}M_{i}$, where $M_{i}$ is accessible for agent $i,i=1,\ldots,N$.
@@
 Figure 1. The structural diagram of the Problem.
-Remark 1
+Remark 1.
 The structural diagram of the studied Problem is shown in Figure 1, where agent $i,i=1,2,\ldots,N$ can only access its own $A_{i}$, $B_{i}$ and communicate with its neighbors.
@@
 It is obvious that each agent makes a decision based on its local information $A_{i}$, $B_{i}$, which aligns with the problem studied in this paper.
-III Preliminaries
-III-A Explicit solution by using all the information
+3 Preliminaries
+3.1 Explicit solution by using all the information
 In this part, we present the optimal solution to the Problem by assuming all the information of matrix parameters are available for the agents.
@@
 To this end, we firstly present the definition of the reachability of system (1) as below.
-Definition 1: A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
+Definition 1.
+A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
 Next, we introduce the Riccati differential equation:
@@
 The optimal solution is then given as follows.
-Theorem 1
+Theorem 1.
 If the terminal state $x_{T}$ is reachable from the initial state $x_{0}$ of system (1), then there exists a solution $\lambda^{*}$ to the following equation:
@@
 Then, the optimal controller of minimizing (2) subject to (1) with $x(0)=x_{0}$ and $x(T)=x_{T}$ is given by:
-$u^{*}(t)=R^{-1}B^{\prime}P(t)x^{*}(t)-R^{-1}B^{\prime}\beta(t),$ (16)
+$u^{*}(t)=-R^{-1}B^{\prime}P(t)x^{*}(t)-R^{-1}B^{\prime}\beta(t),$ (16)
 where $\beta(t)$ obeys the following equation:
@@
 with $x^{*}(0)=x_{0}$.
-Proof: The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
-III-B Numerical solution by using all the information
+Proof. The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
+3.2 Numerical solution by using all the information
 From Theorem 1, the design of optimal controller (16) depends on the solution $P(t)$ of the Riccati differential equation (13), which is nonlinear and typically challenging to solve analytically.
@@
 It is then easy to verify the convergence of the iteration (19) as given below.
-Lemma 1
+Lemma 3.1.
 It holds that
@@
 (2)$\lim_{n\rightarrow\infty}P^{n}(t)=P(t)$, where $P(t)$ satisfies (13).
-Proof: The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
+Proof. The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
 Furthermore, it is immediate to obtain the following result.
-Corollary 1
+Corollary 3.2.
 It holds that
@@
 where $Z(t)=A-BR^{-1}B^{\prime}P(t)$, $\Phi(t)$ satisfies (14), and $\Psi(t)=\Phi^{\prime}(t)^{-1}$.
-Proof: The result follows directly from Lemma 1. $\blacksquare$
-From Lemma 1 and Corollary 1, we have obtained the numerical solution to (13) and (14).
+Proof. The result follows directly from Lemma 3.1. $\blacksquare$
+From Lemma 3.1 and Corollary 3.2, we have obtained the numerical solution to (13) and (14).
 Accordingly, the optimal solution (16) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
@@
 6. Let $u_{c}^{*}(t)=-R^{-1}B^{\prime}P^{n}(t)x_{c}^{*}(t)-R^{-1}B^{\prime}\beta_{c}(t)$ and output $u_{c}^{*}(t)$ and $x_{c}^{*}(t)$.
-Theorem 2
+Theorem 2.
 There exist an integer $n$ and a sufficient small constants $\epsilon_{1}>0$, $\epsilon_{2}>0$ such that output $x_{c}^{*}(t)$, $u_{c}^{*}(t)$ derived by Algorithm 1 satisfies
@@
 where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (16) and (18) by assuming all the information of system parameters are available to make decision.
-Proof: From Lemma 1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$.
+Proof.
+From Lemma 3.1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$.
 This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
@@
 $\blacksquare$
-Remark 2
+Remark 3.3.
 In Algorithm 1, solving the Riccati equation (19) requires $Z^{n}(t)$ and $V^{n}(t)$, which depends on the global system matrices $A$ and $B$.
@@
 This motivates the study of the distributed algorithm.
-IV Main results
+4 Main results
 In this section, we will present a distributed algorithm to derive optimal controller when only partial information is accessible to agent $i$ as provided in Assumptions 1-4 which is divided into three steps.
@@
 where $n=1,2,\cdots$, $k=1,2,\cdots$, $P_{i,0}^{n}(t)=0$, $P_{i.k}^{n}(T)=0$,
-$Z_{i,k}^{n}(t)$ $=$ $Z_{i,k-1}^{n}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$ (31)
-$-Z_{i,k-1}^{n}(t)]+\sum_{j\in N_{i}}[Z_{j,k-1}^{n}(t)-Z_{i,k-1}^{n}(t)],$
-$V_{i,k}^{n}(t)$ $=$ $V_{i,k-1}^{n}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{n-1}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$ (32)
-$-V_{i,k-1}^{n}(t)]+\sum_{j\in N_{i}}[V_{j,k-1}^{n}(t)-V_{i,k-1}^{n}(t)],$
-while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
+$Z_{i,k}^{n}(t)=$ $Z_{i,k-1}^{n}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)-Z_{i,k-1}^{n}(t)]$
+$+\frac{1}{\gamma}\sum_{j\in N_{i}}[Z_{j,k-1}^{n}(t)-Z_{i,k-1}^{n}(t)],$ (31)
+$V_{i,k}^{n}(t)=$ $V_{i,k-1}^{n}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{n-1}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)$
+$-V_{i,k-1}^{n}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[V_{j,k-1}^{n}(t)$
+$-V_{i,k-1}^{n}(t)],$ (32)
+while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, $\gamma>0$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
 By solving (30), the explicit solution of $P_{i,k}^{n}(t)$ is
@@
 It is clear that the convergence of distributed iteration (30) depends on the convergence of distributed iteration (31)-(32) which is shown in the following result.
-Lemma 2
+Lemma 4.1.
 It holds that
@@
 $\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|V_{i,k}^{n}(t)-V(t)\|$ $=$ $0$ (37)
-Proof: The detailed proof is given in Appendix A. $\blacksquare$
+Proof. The detailed proof is given in Appendix A. $\blacksquare$
 Then the following convergence result follows for the distributed iteration (30), (34) and (35).
-Corollary 2
+Corollary 4.2.
 It holds that
@@
 $\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Psi_{i,k}^{n}(t)-\Psi(t)\|$ $=$ $0$ (40)
-Proof: The result follows directly from Lemma 2. $\blacksquare$
+Proof. The result follows directly from Lemma 4.1. $\blacksquare$
 For the convenience of marking, we make the denotations:
@@
 $W_{i,k}$ $=$ $W_{i,k-1}+\alpha_{k}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,k-1}]$ (47)
-$+\sum_{j\in N_{i}}[W_{j,k-1}-W_{i,k-1}],$
+$+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,k-1}-W_{i,k-1}],$
 whose convergence is given below.
-Lemma 3
+Lemma 4.3.
 It holds that
 $W_{i}^{\infty}\triangleq\displaystyle\lim_{k\rightarrow\infty}W_{i,k}=W.$ (48)
-Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
+Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
 Recalling (15) where $\lambda^{*}$ depends not only on the global $W$ but also on the global initial state $x(0)$ and terminal state $x(T)$, we then propose the distributed iteration of $\lambda^{*}$ by
 $\lambda_{i,k}^{*}$ $=$ $\lambda_{i,k-1}^{*}+\alpha_{k}\{(\varrho_{i}^{\infty})^{{\dagger}}[\Phi^{\infty}_{i}(T,0)x_{i,0}-x_{i,T}]$ (49)
-$-\lambda_{i,k-1}^{*}\}+\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
+$-\lambda_{i,k-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
 where
@@
 and show the convergence in the following result.
-Lemma 4
+Lemma 4.4.
 It holds that
 $\lambda_{i}^{\infty,*}\triangleq\displaystyle\lim_{k\rightarrow\infty}\lambda_{i,k}^{*}=\lambda^{*}.$ (51)
-Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
+Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
 Accordingly, the distributed solution for $\beta(t)$ is defined as follows:
@@
 ×$\times{W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\beta_{i}^{\infty}(s)ds-x_{i,k-1}^{*}(t)]$
-$+\sum_{j\in N_{i}}[x_{j,k-1}^{*}(t)-x_{i,k-1}^{*}(t)],$
+$+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,k-1}^{*}(t)-x_{i,k-1}^{*}(t)],$
 and derive the following convergence result.
-Lemma 5
+Lemma 4.5.
 It holds that
 $\displaystyle\lim_{k\rightarrow\infty}x_{i,k}^{*}(t)=x^{*}(t).$ (55)
-Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
-From Lemmas 2-5 and Corollary 2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
+Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
+From Lemmas 4.1-4.5 and Corollary 4.2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
+Theorem 3.
+There exist integers $n,k,\varpi,q,w$ and sufficiently small constants $\psi_{1}>0$, $\psi_{2}>0$ such that output $x_{i,w}^{*}$ and $u_{i}^{*}(t)$ derived by Algorithm 2 satisfy:
+$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
+$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
+where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (16) and (18) by assuming all the information of system parameters are available to make decision.
+Proof.
+From Lemmas 3.1-4.1 and Corollary 4.2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$.
+Together with Lemmas 4.3-4.4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
+$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
+$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
+$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
+Accordingly, by using Lemma 4.5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57).
+This completes the proof.
+$\blacksquare$
+In the following, we generalize the above results to the standard LQ problem without terminal state constraint, i.e. minimizing (2) s. t. (1) with $x(0)=x_{0}$.
+In this case, the optimal controller $u^{\diamond}(t)$ is
+$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
+where $P(t)$ is the solution of equation (13), $x^{\diamond}(t)$ is the solution of the following equation:
+$\dot{x}^{\diamond}(t)$ $=$ $[A-BR^{-1}B^{\prime}P(t)]x^{\diamond}(t),~{}x^{\diamond}(0)=x_{0}.$ (59)
 Algorithm 2 Distributed Algorithm
@@
     $W_{i,\varpi}$ $=$ $W_{i,\varpi-1}+\alpha_{\varpi}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,\varpi-1}]$
-    $+\sum_{j\in N_{i}}[W_{j,\varpi-1}-W_{i,\varpi-1}].$
+    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,\varpi-1}-W_{i,\varpi-1}].$
     • If $\|W_{i,\varpi}-W_{i,\varpi-1}\|<\varsigma$, end the iteration, and derive $W_{i,\varpi}$.
@@
     ×$\times\Psi_{i,k}^{n}(s)ds\Big{]}^{{\dagger}}[\Phi_{i,k}^{n}(T,0)x_{i,0}-x_{i,T}]$
-    $-\lambda_{i,q-1}^{*}\}+\sum_{j\in N_{i}}[\lambda_{j,q-1}^{*}-\lambda_{i,q-1}^{*}].$
+    $-\lambda_{i,q-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,q-1}^{*}-\lambda_{i,q-1}^{*}].$
     • If $\|\lambda_{i,q}^{*}-\lambda_{i,q-1}^{*}\|<\varsigma$, end the iteration, and derive $\lambda_{i,q}^{*}$.
@@
     ×$\times W_{i,\varpi}^{\prime}W_{i,\varpi}\beta_{i}(s)ds-x_{i,w-1}^{*}(t)]$
-    $+\sum_{j\in N_{i}}[x_{j,w-1}^{*}(t)-x_{i,w-1}^{*}(t)].$
+    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,w-1}^{*}(t)-x_{i,w-1}^{*}(t)].$
     • If $\|x_{i,w}^{*}(t)-x_{i,w-1}^{*}(t)\|<\varsigma$, end the iteration, and derive $x_{i,w}^{*}(t)$.
@@
     Let $u_{i}^{*}(t)=-R^{-1/2}W_{i,\varpi}^{\prime}\big{[}P_{i,k}^{n}(t)x_{i,w}^{*}(t)+\beta_{i}(t)\big{]}$ and output $u_{i}^{*}(t)$ and $x_{i,w}^{*}(t)$.
-Theorem 3
-There exist integers $n,k,\varpi,q,w$ and sufficiently small constants $\psi_{1}>0$, $\psi_{2}>0$ such that output $x_{i,w}^{*}$ and $u_{i}^{*}(t)$ derived by Algorithm 2 satisfy:
-$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
-$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
-where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (16) and (18) by assuming all the information of system parameters are available to make decision.
-Proof.
-From Lemmas 1-2 and Corollary 2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$.
-Together with Lemmas 3-4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
-$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
-$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
-$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
-Accordingly, by using Lemma 5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57).
-This completes the proof.
-$\blacksquare$
-In the following, we generalize the above results to the standard LQ problem without terminal state constraint, i.e. minimizing (2) s. t. (1) with $x(0)=x_{0}$.
-In this case, the optimal controller $u^{\diamond}(t)$ is
-$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
-where $P(t)$ is the solution of equation (13), $x^{\diamond}(t)$ is the solution of the following equation:
-$\dot{x}^{\diamond}(t)$ $=$ $[A-BR^{-1}B^{\prime}P(t)]x^{\diamond}(t),~{}x^{\diamond}(0)=x_{0}.$ (59)
 Based on Algorithm 2, the distributed design of the optimal controller (58) is shown in Algorithm 3, the convergence of which is given below.
@@
     $W_{i,\bar{\varpi}}$ $=$ $W_{i,\bar{\varpi}-1}+\alpha_{\bar{\varpi}}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,\bar{\varpi}-1}]$
-    $+\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
+    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
     • If $\|W_{i,\bar{\varpi}}-W_{i,\bar{\varpi}-1}\|<\bar{\varsigma}$, end the iteration, and derive $W_{i,\bar{\varpi}}$.
@@
     $x_{i,\bar{w}}^{\diamond}(t)$ $=$ $x_{i,\bar{w}-1}^{\diamond}(t)+\alpha_{\bar{w}}[\Phi_{i,\bar{k}}^{n}(t,0)x_{i}(0)-x_{i,\bar{w}-1}^{\diamond}(t)]$
-    $+\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
+    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
     • If $\|x_{i,\bar{w}}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)\|<\bar{\varsigma}$, end the iteration, and derive $x_{i,\bar{w}}^{\diamond}(t)$.
@@
     Let $u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(t)x_{i,\bar{w}}^{\diamond}(t)$ and output $u_{i}^{\diamond}(t)$ and $x_{i,\bar{w}}^{\diamond}(t)$.
-Corollary 3
+Corollary 4.6.
 There exist integers $\bar{n},\bar{k},\bar{\varpi},\bar{w}$ and sufficiently small constants $\bar{\psi}_{1}>0$, $\bar{\psi}_{2}>0$ such that output $x_{i,\bar{w}}^{\diamond}(t)$ and $u_{i}^{\diamond}(t)$ derived by Algorithm 3 satisfies
@@
 Proof. The proof is similar to that of Theorem 3 and is thus omitted here.
-Remark 3
+Remark 4.7.
 The distributed solving of the finite-horizon solution (58) has been given in Algorithms 3. When the horizon $T$ is sufficiently large, the output of Algorithm 3
@@
 is an approximation solution to the standard infinite-horizon LQ problem under the assumptions of the stabilization of $(A,B)$ and observability of $(A,Q)$.
-V Simulation Example
+5 Simulation Example
 In this section, two examples are provided to show the effectiveness of the proposed Algorithms.
@@
 $x_{4}(T)=\left[\begin{array}[]{cc}3\\ 6\\ 1\\ 3\end{array}\right],x_{5}(T)=\left[\begin{array}[]{cc}4\\ 7\\ 2\\ 1\end{array}\right].$
-Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8.
+Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$ and $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8.
 It is obvious that all the agents achieve consensus and converge to the centralized solutions from Figure 7-Figure 8.
@@
 $x_{4,0}=\left[\begin{array}[]{cc}5.5\\ 4\end{array}\right],x_{i,T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right],i=1,\cdots,4.$
-By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16.
+By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16.
 It is obvious that the distributed solutions converge to the centralized solutions Figure 15-Figure 16.
@@
 Figure 16. The trajectories of $u_{i}^{*}(t)$.
-VI Conclusions
+6 Conclusions
 This paper studied the LQ optimal control problem for continuous-time systems with terminal state constraints.
@@
 Appendix A
-Proof of Lemma 2
 We will prove the result by using mathematical induction. For $n=1$, it is obvious from (31)-(32) that
 $Z_{i,k}^{1}(t)$ $=$ $Z_{i,k-1}^{1}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$ (86)
-$-Z_{i,k-1}^{1}(t)]+\sum_{j\in N_{i}}[Z_{j,k}^{1}(t)-Z_{i,k-1}^{1}(t)],$
+$-Z_{i,k-1}^{1}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[Z_{j,k-1}^{1}(t)-Z_{i,k-1}^{1}(t)],$
 $V_{i,k}^{1}(t)$ $=$ $V_{i,k-1}^{1}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{0}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$ (87)
-$-V_{i,k-1}^{1}(t)]+\sum_{j\in N_{i}}[V_{j,k}^{1}(t)-V_{i,k-1}^{1}(t)].$
+$-V_{i,k-1}^{1}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[V_{j,k-1}^{1}(t)-V_{i,k-1}^{1}(t)].$
 It is aimed to show
@@
 $\mathcal{V}_{k}^{1}(t)$ $=$ $(\mathcal{A}\otimes I_{n})\mathcal{V}_{k-1}^{1}(t)+\alpha_{k}\Theta_{k-1}^{1}(t),$
-where $\mathcal{A}=I_{N}-\mathcal{L}$, $\mathcal{L}$ is the Laplacian matrix [31].
+where $\mathcal{A}=I_{N}-\frac{1}{\gamma}\mathcal{L}$, $\mathcal{L}$ is the Laplacian matrix [31].
 It is obvious that $\mathcal{Z}_{k}^{1}(t)$ and $\mathcal{V}_{k}^{1}(t)$ are bounded.
@@
 According to (86)-(87), we have that
-$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
+$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
 $-\alpha_{k}(U^{\prime}\otimes I_{n})\Pi_{k-1}^{1}(t),$
-$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
+$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
 $-\alpha_{k}(U^{\prime}\otimes I_{n})\Theta_{k-1}^{1}(t).$
@@
 ×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|.$
-In view of the fact that the graph is connected, it yields that
+Since the graph is connected, there exists a constant $\gamma>0$ such that the following inequality holds
 $\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\leq c\rho^{k},$

### v4

@@
 The main contribution is to present a distributed algorithm to derive the optimal controller which is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state.
-The effectiveness of the proposed algorithm is verified by two numerical examples.
-keywords:
+Furthermore, the proposed distributed iteration method is extended to solve the consensus control problem for heterogeneous multi-agent systems, achieving the globally optimal performance of the system.
+The effectiveness of the proposed algorithm is verified by two numerical examples, where the performance index under the proposed distributed controller is smaller than that under the commonly used consensus control.
+Index Terms:
 LQ optimal control, Terminal state constraint, Distributed algorithm.
-thanks: Corresponding author
-1 Introduction
+I Introduction
 Linear quadratic (LQ) optimal control is one of the core fundamental problem in modern control theory which has wide applications in plenty of fields including power system scheduling [1], economics [2] and autonomous driving [3].
@@
 [28] discussed the distributed consensus optimal control strategy for multi-agent systems with a general directed graph.
-[29] investigated a globally optimal distributed formation control algorithm.
+[29] studied the design of suboptimal distributed control for the identical decoupled dynamical systems.
 [30] considered the online distributed convex optimization problem with local dynamic loss function sums under unbalanced graphs.
@@
 Specifically, the distributed algorithm is consisting of the distributed iterations for the Riccati equation, a backward differential equation driven by the optimal Lagrange multiplier and the optimal state.
+Furthermore, we solve the optimal consensus control problem using the proposed distributed iteration method.
+Numerical examples demonstrate the effectiveness of the proposed algorithms, where the performance index under the proposed distributed controller is smaller than that under the commonly used consensus control.
 The remainder of the paper is organized as follows.
@@
 Section IV gives the distributed algorithm for the optimal controller.
-Numerical example is given in Section V.
+Section V gives the distributed solution to the optimal consensus control problem.
+Numerical examples are given in Section VI.
 The conclusion is outlined in the last Section.
 Detailed proofs of the results are given in Appendix.
-2 Problem formulation
+The following notations will be used throughout this paper: $R^{n}$ denotes the set of $n$-dimensional vectors; $x^{\prime}$ denotes the transpose of $x$.
+II is identity matrices with compatible dimensions.
+II Problem formulation
 Consider the following continuous-time system:
 $\dot{x}(t)=Ax(t)+Bu(t),$ (1)
-where $x\in R^{n}$ is the state, $u\in R^{m}$ is the control input.
+where $x(t)\in R^{n}$ is the state, $u(t)\in R^{m}$ is the control input.
 $A$ and $B$ are constant matrices with compatible dimensions.
@@
 The cost function is given by
-[…]
+$J=$ $~\int_{0}^{T}[x^{\prime}(t)Qx(t)+u^{\prime}(t)Ru(t)]dt,$ (2)
 where $Q$ is a positive semi-definite matrix with compatible dimension, and $R$ is a positive-definite matrix with compatible dimension.
@@
 In details, the following assumptions are given to show the detailed information available to agent $i$.
-Assumption 1.
+Assumption 1
 There exist matrices $A_{i},B_{i},i=1,\ldots,N$ such that $A=\frac{1}{N}\sum_{i=1}^{N}A_{i}$, $B=\frac{1}{N}\sum_{i=1}^{N}B_{i}$.
 In particular, $A_{i}$ and $B_{i}$ are accessible for agent $i,i=1,\ldots,N$.
-Assumption 2.
+Assumption 2
 There exist matrices $Q_{i},i=1,\ldots,N$ such that $Q=\frac{1}{N}\sum_{i=1}^{N}Q_{i}$, and $Q_{i}$ is accessible for agent $i,i=1,\ldots,N$. The matrix $R$ is accessible to each agent.
-Assumption 3.
+Assumption 3
 There exist $x_{i,0},x_{i,T},i=1,\ldots,N$ such that $x(0)=x_{0}=\frac{1}{N}\sum_{i=1}^{N}x_{i,0}$ and $x(T)=x_{T}=\frac{1}{N}\sum_{i=1}^{N}x_{i,T}$.
 In particular, $x_{i,0},x_{i,T}$ accessible to agent $i,i=1,\ldots,N$.
-Assumption 4.
-There exists matrices $M_{i},i=1,\ldots,N$ such that $BR^{-1}B^{\prime}=\frac{1}{N}\sum_{i=1}^{N}M_{i}^{\prime}M_{i}$, where $M_{i}$ is accessible for agent $i,i=1,\ldots,N$.
+Assumption 4
+There exists matrices $M_{i},i=1,\ldots,N$ such that $BR^{-1}B^{\prime}=\frac{1}{N}\sum_{i=1}^{N}M_{i}M_{i}^{\prime}$, where $M_{i}$ is accessible for agent $i,i=1,\ldots,N$.
 Based on the above description, the studied problem is concluded as follows.
-Problem.
+Problem 1.
 The aim of the paper is to design a distributed algorithm for the optimal controller which minimizes the cost function (2) subject to system (1) and ensures the state to satisfy $x(0)=x_{0}$ and $x(T)=x_{T}$ for each agent $i$ under Assumptions 1-4.
-Figure 1. The structural diagram of the Problem.
-Remark 1.
-The structural diagram of the studied Problem is shown in Figure 1, where agent $i,i=1,2,\ldots,N$ can only access its own $A_{i}$, $B_{i}$ and communicate with its neighbors.
-The problem has important applications in the fields of urban intelligent surveillance, disaster emergency response, and collaborative search and rescue.
-For example, in the distributed control problem of heterogeneous multi-agent systems (HMASs) consisting of unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs) [32], the dynamics of the $i$ agent can be described by
-$\dot{x}_{i}(t)=\mathcal{A}_{i}x_{i}(t)+\mathcal{B}_{i}u_{i}(t),i=1,..,N.$ (3)
-where $x_{i}(t)\in R^{n}$, $u_{i}(t)\in R^{m}$.
-By defining $x(t)=[x_{1}(t)^{\prime}~{}\cdots~{}x_{N}(t)^{\prime}]^{\prime}$, $u(t)=[u_{1}(t)^{\prime}~{}\cdots~{}u_{N}(t)^{\prime}]^{\prime}$, the system (3) can be expressed in the global form of system (1) with $A=\frac{1}{N}\sum_{i=1}^{N}A_{i}$, $B=\frac{1}{N}\sum_{i=1}^{N}B_{i}$, where
-$A_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&\mathcal{A}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],B_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&\mathcal{B}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right].$ (12)
-It is obvious that each agent makes a decision based on its local information $A_{i}$, $B_{i}$, which aligns with the problem studied in this paper.
-3 Preliminaries
-3.1 Explicit solution by using all the information
-In this part, we present the optimal solution to the Problem by assuming all the information of matrix parameters are available for the agents.
+Remark 1
+It is noted that Assumptions 1-4 are meaningful and have practical applications in network systems [31] and multi-agent systems, where the application of multi-agent systems is studied in Section V.
+In addition, we also note that although the system matrix was assumed to have a decomposable form under Assumptions 1-4, solving Problem 1 cannot be simply decomposed into the sum of $N$ subproblems, for example, in distributed optimal control problem where Assumptions 1-4 hold naturally, the derivation of the optimal solution is NP-hard [29].
+III Preliminaries
+III-A Explicit solution by using all the information
+In this part, we present the optimal solution to the Problem 1 by assuming all the information of matrix parameters are available for the agents.
 Noting that the terminal state is predetermined, thus it is necessary to ensure that the system can be driven from the initial state to the terminal state.
 To this end, we firstly present the definition of the reachability of system (1) as below.
-Definition 1.
-A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
+Definition 1: A state $x_{T}$ is said to be reachable at time $T$ from the initial state $x_{0}$ of system (1) if there exists a controller $u(t)$, $t\in[0,T]$ such that the solution $x(t)$ of the system (1) satisfies $x(T)=x_{T}$.
 Next, we introduce the Riccati differential equation:
@@
 The optimal solution is then given as follows.
-Theorem 1.
+Theorem 1
 If the terminal state $x_{T}$ is reachable from the initial state $x_{0}$ of system (1), then there exists a solution $\lambda^{*}$ to the following equation:
-[…]
+$\Big[\int_{0}^{T}\Phi(T,s)BR^{-1}B^{\prime}\Psi(s,T)ds\Big]\lambda=\Phi(T,0)x_{0}-x_{T}.$ (5)
 Then, the optimal controller of minimizing (2) subject to (1) with $x(0)=x_{0}$ and $x(T)=x_{T}$ is given by:
@@
 where $\beta(t)$ obeys the following equation:
-$\dot{\beta}(t)=-[A^{\prime}-P(t)BR^{-1}B^{\prime}]\beta(t),~{}\beta(T)=\lambda^{*}$ (17)
+$\dot{\beta}(t)=-[A^{\prime}-P(t)BR^{-1}B^{\prime}]\beta(t),~\beta(T)=\lambda^{*},$ (7)
 and $x^{*}(t)$ obeys the following equation:
@@
 with $x^{*}(0)=x_{0}$.
-Proof. The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
-3.2 Numerical solution by using all the information
+Proof: The proof follows directly from Theorem 1 in [33] and is thus omitted here. $\blacksquare$
+III-B Numerical solution by using all the information
 From Theorem 1, the design of optimal controller (6) depends on the solution $P(t)$ of the Riccati differential equation (3), which is nonlinear and typically challenging to solve analytically.
@@
 with $P^{n}(T)=0$, $P^{0}(t)=0$, and give its solution by
-[…]
+$P^{n}(t)=\Psi^{n}(t)\big[\int_{t}^{T}\Phi^{n}(s)^{\prime}V^{n}(s)\Phi^{n}(s)ds\big]\Psi^{n}(t)^{\prime},$ (12)
 where $\Phi^{n}(t)$ and $\Psi^{n}(t)$ satisfy the following iteration respectively:
@@
 It is then easy to verify the convergence of the iteration (9) as given below.
-Lemma 3.1.
+Lemma 1
 It holds that
 (1) $P^{n}(t)\geq P^{n+1}(t)$;
-(2)$\lim_{n\rightarrow\infty}P^{n}(t)=P(t)$, where $P(t)$ satisfies (13).
-Proof. The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
+[…]
+Proof: The proof can be directly obtained from Theorem 7.2 in [34] and is thus omitted here. $\blacksquare$
 Furthermore, it is immediate to obtain the following result.
-Corollary 3.2.
+Corollary 1
 It holds that
@@
 where $Z(t)=A-BR^{-1}B^{\prime}P(t)$, $\Phi(t)$ satisfies (4), and $\Psi(t)=\Phi^{\prime}(t)^{-1}$.
-Proof. The result follows directly from Lemma 3.1. $\blacksquare$
-From Lemma 3.1 and Corollary 3.2, we have obtained the numerical solution to (13) and (14).
+Proof: The result follows directly from Lemma 1. $\blacksquare$
+From Lemma 1 and Corollary 1, we have obtained the numerical solution to (3) and (4).
 Accordingly, the optimal solution (6) can be numerically obtained, and then the detailed algorithm is given in Algorithm 1.
 Algorithm 1 Centralized Algorithm
-Input: Let the initial matrices be $P^{n}(T)=0,n=1,2,3,\ldots$, $P^{0}(t)=0,t\in[0,T]$, and select a sufficient small constant $\varepsilon>0$.
+Input: Let the initial matrices be $P^{n}(T)=0,n=1,2,3,\ldots$, $P^{0}(t)=0,t\in[0,T]$, and select a sufficiently small constant $\varepsilon>0$.
 Output: The optimal states $x_{c}^{*}(t)$ and controller $u^{*}_{c}(t)$.
@@
 3. Derive $\lambda_{c}^{*}$ by using
-[…]
+$\lambda_{c}^{*}=\Big[\int_{0}^{T}\Phi^{n}(T,s)BR^{-1}B^{\prime}\Psi^{n}(s,T)ds\Big]^{{\dagger}}[\Phi^{n}(T,0)x_{0}-x_{T}].$
 4. Solve the equation
-$\dot{\beta}_{c}(t)=-Z^{n}(t)\beta_{c}(t),\beta_{c}(T)=\lambda_{c}^{*},$
+$\dot{\beta}_{c}(t)=-Z^{n}(t)^{\prime}\beta_{c}(t),\beta_{c}(T)=\lambda_{c}^{*},$
 and obtain $\beta_{c}(t)$.
@@
 6. Let $u_{c}^{*}(t)=-R^{-1}B^{\prime}P^{n}(t)x_{c}^{*}(t)-R^{-1}B^{\prime}\beta_{c}(t)$ and output $u_{c}^{*}(t)$ and $x_{c}^{*}(t)$.
-Theorem 2.
+Theorem 2
 There exist an integer $n$ and a sufficient small constants $\epsilon_{1}>0$, $\epsilon_{2}>0$ such that output $x_{c}^{*}(t)$, $u_{c}^{*}(t)$ derived by Algorithm 1 satisfies
@@
 where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (6) and (8) by assuming all the information of system parameters are available to make decision.
-Proof.
-From Lemma 3.1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$.
+Proof: From Lemma 1, there always exist an integer $n$ and a sufficient small constant $\varepsilon>0$ such that $\|P^{n}(t)-P(t)\|\leq\varepsilon,t\in[0,T]$.
 This further implies that there exist sufficient small constants $\varepsilon_{1}>0,\varepsilon_{2}>0$ such that
 $\|\lambda_{c}-\lambda^{*}\|\leq\varepsilon_{1},$
-$\|\beta_{c}(t)-\beta(t)\|\leq\varepsilon_{2},$
+$\|\beta_{c}(t)-\beta(t)\|\leq\varepsilon_{2}.$
 Accordingly, there exists a sufficient small constant $\epsilon_{1}>0$, $\epsilon_{2}>0$ such that $x_{c}^{*}(t)$ and $u_{c}^{*}(t)$ obtained by running Algorithm 1 satisfies (18) and (19).
@@
 $\blacksquare$
-Remark 3.3.
+Remark 2
 In Algorithm 1, solving the Riccati equation (9) requires $Z^{n}(t)$ and $V^{n}(t)$, which depends on the global system matrices $A$ and $B$.
@@
 This motivates the study of the distributed algorithm.
-4 Main results
+IV Main results
 In this section, we will present a distributed algorithm to derive optimal controller when only partial information is accessible to agent $i$ as provided in Assumptions 1-4 which is divided into three steps.
@@
 $\dot{P}_{i,k}^{n}(t)$ $=$ $-Z_{i,k}^{n}(t)^{\prime}P_{i,k}^{n}(t)-P_{i,k}^{n}(t)Z_{i,k}^{n}(t)-V_{i,k}^{n}(t),$ (20)
-where $n=1,2,\cdots$, $k=1,2,\cdots$, $P_{i,0}^{n}(t)=0$, $P_{i.k}^{n}(T)=0$,
+where $n=1,2,\cdots$, $k=1,2,\cdots$, $P_{i,k}^{0}(t)=0$, $P_{i,0}^{n}(t)=0$, $P_{i,k}^{n}(T)=0$,
 $Z_{i,k}^{n}(t)=$ $Z_{i,k-1}^{n}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{n-1}(t)-Z_{i,k-1}^{n}(t)]$
@@
 $-V_{i,k-1}^{n}(t)],$ (22)
-while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}\leq\infty$, $\gamma>0$, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
+while $\alpha_{k}$ is a step size satisfying $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, $\gamma>0$ is selected such that the matrix $I_{N}-\frac{1}{\gamma}\mathcal{L}-\frac{1}{N}1_{N}1_{N}^{\prime}$ is Hurwitz, ensuring the convergence of the distributed iterations, and $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$.
 By solving (20), the explicit solution of $P_{i,k}^{n}(t)$ is
 $P_{i,k}^{n}(t)$ $=$ $\Psi_{i,k}^{n}(t)\big[\int_{t}^{T}\Phi_{i,k}^{n}(s)^{\prime}V_{i,k}^{n}(s)$ (23)
-×$\times\Phi_{i,k}^{n}(s)ds\big{]}\Psi_{i,k}^{n}(t)^{\prime}.$
+$\times\Phi_{i,k}^{n}(s)ds\big]\Psi_{i,k}^{n}(t)^{\prime},$
 where $\Phi_{i,k}^{n}(t)$ and $\Psi_{i,k}^{n}(t)$ satisfy the following iteration:
@@
 It is clear that the convergence of distributed iteration (20) depends on the convergence of distributed iteration (21)-(22) which is shown in the following result.
-Lemma 4.1.
+Lemma 2
 It holds that
 $\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|Z_{i,k}^{n}(t)-Z(t)\|$ $=$ $0,$ (26)
-$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|V_{i,k}^{n}(t)-V(t)\|$ $=$ $0$ (37)
-Proof. The detailed proof is given in Appendix A. $\blacksquare$
+$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|V_{i,k}^{n}(t)-V(t)\|$ $=$ $0.$ (27)
+Proof: The detailed proof is given in Appendix A. $\blacksquare$
 Then the following convergence result follows for the distributed iteration (20), (24) and (25).
-Corollary 4.2.
+Corollary 2
 It holds that
@@
 $\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Phi_{i,k}^{n}(t)-\Phi(t)\|$ $=$ $0,$ (29)
-$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Psi_{i,k}^{n}(t)-\Psi(t)\|$ $=$ $0$ (40)
-Proof. The result follows directly from Lemma 4.1. $\blacksquare$
+$\displaystyle\lim_{n\rightarrow\infty}\lim_{k\rightarrow\infty}\|\Psi_{i,k}^{n}(t)-\Psi(t)\|$ $=$ $0.$ (30)
+Proof: The result follows directly from Lemma 2. $\blacksquare$
 For the convenience of marking, we make the denotations:
@@
 Secondly, to derive the distributed solution of $\lambda^{*}$, we denote
-[…]
+$\varrho=\int_{0}^{T}\Phi(s)W^{\prime}W\Psi(s)ds,$ (36)
 where $W=R^{-\frac{1}{2}}B^{\prime}$ and introduce the following distributed iteration for matrix $W$ as:
@@
 $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,k-1}-W_{i,k-1}],$
-whose convergence is given below.
-Lemma 4.3.
+with $W_{i,0}=0$, whose convergence is given below.
+Lemma 3
 It holds that
 $W_{i}^{\infty}\triangleq\displaystyle\lim_{k\rightarrow\infty}W_{i,k}=W.$ (38)
-Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
+Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
 Recalling (5) where $\lambda^{*}$ depends not only on the global $W$ but also on the global initial state $x(0)$ and terminal state $x(T)$, we then propose the distributed iteration of $\lambda^{*}$ by
 $\lambda_{i,k}^{*}$ $=$ $\lambda_{i,k-1}^{*}+\alpha_{k}\{(\varrho_{i}^{\infty})^{{\dagger}}[\Phi^{\infty}_{i}(T,0)x_{i,0}-x_{i,T}]$ (39)
-$-\lambda_{i,k-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}].$
-where
-[…]
-and show the convergence in the following result.
-Lemma 4.4.
+$-\lambda_{i,k-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,k-1}^{*}-\lambda_{i,k-1}^{*}],$
+with $\lambda_{i,0}^{*}=0$ and
+$\varrho_{i}^{\infty}=\int_{0}^{T}\Phi^{\infty}_{i}(s){W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\Psi^{\infty}_{i}(s)ds.$ (40)
+The convergence of this iteration in the following result.
+Lemma 4
 It holds that
 $\lambda_{i}^{\infty,*}\triangleq\displaystyle\lim_{k\rightarrow\infty}\lambda_{i,k}^{*}=\lambda^{*}.$ (41)
-Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
+Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
 Accordingly, the distributed solution for $\beta(t)$ is defined as follows:
-$\dot{\beta}_{i}^{\infty}(t)$ $=$ $-Z^{\infty}_{i}(t)\beta_{i}^{\infty}(t),$ (52)
+$\dot{\beta}_{i}^{\infty}(t)$ $=$ $-Z^{\infty}_{i}(t)^{\prime}\beta_{i}^{\infty}(t),$ (42)
 $\beta_{i}^{\infty}(T)$ $=$ $\lambda_{i}^{\infty,*}.$
@@
 Thirdly, we give the distributed solution of $x^{*}(t)$. In fact, by solving (5), the explicit solution of $x^{*}(t)$ is given by
-[…]
+$x^{*}(t)=\Phi(t,0)x_{0}-\int_{0}^{t}\Phi(t,s)W^{\prime}W\beta(s)ds,$ (43)
 which relies on the global initial state $x_{0}$. To avoid the using of global information, we give the distributed iteration as follows:
 $x_{i,k}^{*}(t)$ $=$ $x_{i,k-1}^{*}(t)+\alpha_{k}[\Phi_{i}^{\infty}(t,0)x_{i,0}-\int_{0}^{t}\Phi_{i}^{\infty}(t,s)$ (44)
-×$\times{W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\beta_{i}^{\infty}(s)ds-x_{i,k-1}^{*}(t)]$
+$\times{W_{i}^{\infty}}^{\prime}W_{i}^{\infty}\beta_{i}^{\infty}(s)ds-x_{i,k-1}^{*}(t)]$
 $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,k-1}^{*}(t)-x_{i,k-1}^{*}(t)],$
-and derive the following convergence result.
-Lemma 4.5.
+with $x_{i,0}^{*}(t)=0$ and derive the following convergence result.
+Lemma 5
 It holds that
 $\displaystyle\lim_{k\rightarrow\infty}x_{i,k}^{*}(t)=x^{*}(t).$ (45)
-Proof. The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
-From Lemmas 4.1-4.5 and Corollary 4.2, we have established the distributed solutions to (13) and (14). This enables the distributed design of the optimal controller (16) as shown in Algorithm 2.
-Theorem 3.
-There exist integers $n,k,\varpi,q,w$ and sufficiently small constants $\psi_{1}>0$, $\psi_{2}>0$ such that output $x_{i,w}^{*}$ and $u_{i}^{*}(t)$ derived by Algorithm 2 satisfy:
-$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (56)
-$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (57)
-where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (16) and (18) by assuming all the information of system parameters are available to make decision.
-Proof.
-From Lemmas 3.1-4.1 and Corollary 4.2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$.
-Together with Lemmas 4.3-4.4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
-$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
-$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
-$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4},$
-Accordingly, by using Lemma 4.5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (56) and (57).
-This completes the proof.
-$\blacksquare$
-In the following, we generalize the above results to the standard LQ problem without terminal state constraint, i.e. minimizing (2) s. t. (1) with $x(0)=x_{0}$.
-In this case, the optimal controller $u^{\diamond}(t)$ is
-$u^{\diamond}(t)=-R^{-1}B^{\prime}P(t)x^{\diamond}(t),$ (58)
-where $P(t)$ is the solution of equation (13), $x^{\diamond}(t)$ is the solution of the following equation:
-$\dot{x}^{\diamond}(t)$ $=$ $[A-BR^{-1}B^{\prime}P(t)]x^{\diamond}(t),~{}x^{\diamond}(0)=x_{0}.$ (59)
+Proof: The proof is similar to Lemma 2 and is omitted here. $\blacksquare$
+From Lemmas 2-5 and Corollary 2, we have established the distributed solutions to (3) and (4). This enables the distributed design of the optimal controller (6) as shown in Algorithm 2.
 Algorithm 2 Distributed Algorithm
-Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$, $P_{i,k}^{0}(t)=0$, $W_{i,0}=0$, $\lambda_{i,0}^{*}=0$, and $x_{i,0}^{*}(t)=0$ and select a sufficient small constant $\varsigma>0$.
+Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{n}(t)=0$, $P_{i,k}^{0}(t)=0$, $W_{i,0}=0$, $\lambda_{i,0}^{*}=0$, and $x_{i,0}^{*}(t)=0$ and select a sufficiently small constant $\varsigma>0$.
 Output: The optimal states $x_{i,w}^{*}(t)$ and distributed controller $u_{i}^{*}(t)$, $i=1,2,\ldots,N$.
@@
     $\lambda_{i,q}^{*}$ $=$ $\lambda_{i,q-1}^{*}+\alpha_{q}\bigg\{\Big[\int_{0}^{T}\Phi_{i,k}^{n}(s)W_{i,\varpi}^{\prime}W_{i,\varpi}$
-    ×$\times\Psi_{i,k}^{n}(s)ds\Big{]}^{{\dagger}}[\Phi_{i,k}^{n}(T,0)x_{i,0}-x_{i,T}]$
+    $\times\Psi_{i,k}^{n}(s)ds\Big]^{{\dagger}}[\Phi_{i,k}^{n}(T,0)x_{i,0}-x_{i,T}]$
     $-\lambda_{i,q-1}^{*}\}+\frac{1}{\gamma}\sum_{j\in N_{i}}[\lambda_{j,q-1}^{*}-\lambda_{i,q-1}^{*}].$
@@
     Solve the equation
-    $\dot{\beta}_{i}(t)=-Z_{i,k}^{n}(t)\beta_{i}(t),~{}\beta_{i}(T)=\lambda_{i,q}^{*},$
+    $\dot{\beta}_{i}(t)=-Z_{i,k}^{n}(t)^{\prime}\beta_{i}(t),~\beta_{i}(T)=\lambda_{i,q}^{*},$
     and obtain $\beta_{i}(t)$.
@@
     $x_{i,w}^{*}(t)$ $=$ $x_{i,w-1}^{*}(t)+\alpha_{w}[\Phi_{i,k}^{n}(t,0)x_{i,0}-\int_{0}^{t}\Phi_{i,k}^{n}(t,s)$
-    ×$\times W_{i,\varpi}^{\prime}W_{i,\varpi}\beta_{i}(s)ds-x_{i,w-1}^{*}(t)]$
+    $\times W_{i,\varpi}^{\prime}W_{i,\varpi}\beta_{i}(s)ds-x_{i,w-1}^{*}(t)]$
     $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,w-1}^{*}(t)-x_{i,w-1}^{*}(t)].$
@@
   6. 6.
-    Let $u_{i}^{*}(t)=-R^{-1/2}W_{i,\varpi}^{\prime}\big{[}P_{i,k}^{n}(t)x_{i,w}^{*}(t)+\beta_{i}(t)\big{]}$ and output $u_{i}^{*}(t)$ and $x_{i,w}^{*}(t)$.
-Based on Algorithm 2, the distributed design of the optimal controller (58) is shown in Algorithm 3, the convergence of which is given below.
-Algorithm 3 Distributed Algorithm without Terminal State Constraint
-Input: Let the initial matrices be $Z_{i,0}^{n}(t)=0$, $V_{i,0}^{\bar{n}}(t)=0$, $P_{i,\bar{n}}^{0}(t)=0$, $W_{i,0}=0$, and $x_{i,0}^{\diamond}(t)=0$ and select a sufficient small constant $\bar{\varsigma}>0$.
-Output: The distributed controller $u_{i}^{\diamond}(t)$, $i=1,2,\ldots,N$.
-  1. 1.
-    For $\bar{n}=1,2,\cdots$,
-    • For $\bar{k}=1,2,3,\cdots$, run the iteration (31)-(32) and obtain $Z_{i,\bar{k}}^{\bar{n}}(t)$, $V_{i,\bar{k}}^{\bar{n}}(t)$.
-If $\|Z_{i,\bar{k}}^{n}(t)-Z_{i,\bar{k}-1}^{\bar{n}}(t)\|<\bar{\varsigma}$ and $\|V_{i,\bar{k}}^{n}(t)-V_{i,\bar{k}-1}^{\bar{n}}(t)\|<\bar{\varsigma}$, end the iteration, and derive $Z_{i,\bar{k}}^{\bar{n}}(t)$ and $V_{i,\bar{k}}^{\bar{n}}(t)$.
-    • Solve the equation (30) and obtain $P_{i,\bar{k}}^{\bar{n}}(t)$.
-If $\|P_{i,\bar{k}}^{\bar{n}}(t)-P_{i,\bar{k}}^{\bar{n}-1}(t)\|<\bar{\varsigma}$, end the iteration, and derive $P^{\bar{n}}_{i,\bar{k}}(t)$, $\Phi^{\bar{n}}_{i,\bar{k}}(t)$.
-  2. 2.
-    For $\bar{\varpi}=1,2,\cdots$, run the iteration
-    $W_{i,\bar{\varpi}}$ $=$ $W_{i,\bar{\varpi}-1}+\alpha_{\bar{\varpi}}[R^{-\frac{1}{2}}B_{i}^{\prime}-W_{i,\bar{\varpi}-1}]$
-    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[W_{j,\bar{\varpi}-1}-W_{i,\bar{\varpi}-1}].$
-    • If $\|W_{i,\bar{\varpi}}-W_{i,\bar{\varpi}-1}\|<\bar{\varsigma}$, end the iteration, and derive $W_{i,\bar{\varpi}}$.
-  3. 3.
-    For $\bar{w}=1,2,\cdots$, run the iteration
-    $x_{i,\bar{w}}^{\diamond}(t)$ $=$ $x_{i,\bar{w}-1}^{\diamond}(t)+\alpha_{\bar{w}}[\Phi_{i,\bar{k}}^{n}(t,0)x_{i}(0)-x_{i,\bar{w}-1}^{\diamond}(t)]$
-    $+\frac{1}{\gamma}\sum_{j\in N_{i}}[x_{j,\bar{w}-1}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)].$
-    • If $\|x_{i,\bar{w}}^{\diamond}(t)-x_{i,\bar{w}-1}^{\diamond}(t)\|<\bar{\varsigma}$, end the iteration, and derive $x_{i,\bar{w}}^{\diamond}(t)$.
-  4. 4.
-    Let $u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(t)x_{i,\bar{w}}^{\diamond}(t)$ and output $u_{i}^{\diamond}(t)$ and $x_{i,\bar{w}}^{\diamond}(t)$.
-Corollary 4.6.
-There exist integers $\bar{n},\bar{k},\bar{\varpi},\bar{w}$ and sufficiently small constants $\bar{\psi}_{1}>0$, $\bar{\psi}_{2}>0$ such that output $x_{i,\bar{w}}^{\diamond}(t)$ and $u_{i}^{\diamond}(t)$ derived by Algorithm 3 satisfies
-$\|x_{i,\bar{w}}^{\diamond}(t)-x^{\diamond}(t)\|\leq\bar{\psi}_{1},$ (60)
-$\|u_{i}^{\diamond}(t)-u^{\diamond}(t)\|\leq\bar{\psi}_{2},$ (61)
-where $x^{\diamond}(t)$ and $u^{\diamond}(t)$ are the optimal solution of (59) and (58) to the standard LQ problem without terminal state constraint by assuming all the information of system parameters are available to make decision.
-Proof. The proof is similar to that of Theorem 3 and is thus omitted here.
-Remark 4.7.
-The distributed solving of the finite-horizon solution (58) has been given in Algorithms 3. When the horizon $T$ is sufficiently large, the output of Algorithm 3
-$u_{i}^{\diamond}(t)=-R^{-1/2}W_{i,\bar{\varpi}}^{\prime}P_{i,\bar{k}}^{\bar{n}}(T)x_{i,\bar{w}}^{\diamond}(t).$ (62)
-is an approximation solution to the standard infinite-horizon LQ problem under the assumptions of the stabilization of $(A,B)$ and observability of $(A,Q)$.
-5 Simulation Example
+    Let $u_{i}^{*}(t)=-R^{-1/2}W_{i,\varpi}\big[P_{i,k}^{n}(t)x_{i,w}^{*}(t)+\beta_{i}(t)\big]$ and output $u_{i}^{*}(t)$ and $x_{i,w}^{*}(t)$.
+Theorem 3
+There exist integers $n,k,\varpi,q,w$ and sufficiently small constants $\psi_{1}>0$, $\psi_{2}>0$ such that output $x_{i,w}^{*}$ and $u_{i}^{*}(t)$ derived by Algorithm 2 satisfy:
+$\|x_{i,w}^{*}(t)-x^{*}(t)\|\leq\psi_{1},$ (46)
+$\|u_{i}^{*}(t)-u^{*}(t)\|\leq\psi_{2},$ (47)
+where $x^{*}(t)$ and $u^{*}(t)$ are the optimal solution of equations (6) and (8) by assuming all the information of system parameters are available to make decision.
+Proof.
+From Lemmas 1-2 and Corollary 2, there always exists integer $n$, $k$ and a sufficient small constant $\sigma_{1}>0$ such that $\|P_{i,k}^{n}(t)-P(t)\|\leq\sigma_{1},t\in[0,T]$.
+Together with Lemmas 3-4, this further implies that there exist iteration numbers $\varpi,q$ and sufficient small constants $\sigma_{i}>0,i=2,\cdots,4$ such that
+$\|W_{i,\varpi}-W\|\leq\sigma_{2},$
+$\|\lambda_{i,q}^{*}-\lambda^{*}\|\leq\sigma_{3},$
+$\|\beta_{i}(t)-\beta(t)\|\leq\sigma_{4}.$
+Accordingly, by using Lemma 5, there exist iteration numbers $w$ and sufficient small constants $\psi_{1}>0$, $\psi_{2}>0$ such that the outputs $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$ obtained by running Algorithm 2 satisfies (49) and (47).
+This completes the proof.
+$\blacksquare$
+V Application to the Optimal Consensus Control Problem
+In this section, we will utilize Algorithm 2 to solve the consensus problem in multi-agent systems, which has broad applications in fields such as unmanned aerial vehicle formation, sensor networks, smart grid dispatch, and distributed robotic collaboration.
+Specifically, consider the following heterogeneous multi-agent system consisting of $N$ agents with dynamics described by:
+$\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),$ (48)
+where $x_{i}(t)\in R^{n}$, $u_{i}(t)\in R^{m}$, $i=1,..,N$. The initial state is given by $x_{i}(0)=x_{i,0}$.
+To apply Algorithm 2 and achieve consensus, we define the following performance index:
+$J$ $=\int_{0}^{\infty}\Big[\sum_{i=1}^{N}\sum_{j\in N_{i}}(x_{i}(t)-x_{j}(t))^{\prime}Q_{ij}(x_{i}(t)-x_{j}(t))$
+$\quad\quad\quad~~+u^{\prime}(t)Ru(t)\Big]dt$
+$=\int_{0}^{\infty}[x^{\prime}(t)\tilde{Q}x(t)+u^{\prime}(t)Ru(t)]dt,$ (49)
+where $x(t)=[x_{1}^{\prime}(t)~\cdots~x_{N}^{\prime}(t)]^{\prime}$, $u(t)=[u_{1}^{\prime}(t)~\cdots~u_{N}^{\prime}(t)]^{\prime}$, $R=diag\{R_{1},R_{2},\ldots,R_{N}\}$ is a positive-definite matrix with compatible dimension, $Q_{ij}$, $j\in N_{i}$ is a positive definite matrix with compatible dimension, and
+$\tilde{Q}$ $=\begin{bmatrix}\tilde{Q}_{11}&\tilde{Q}_{12}&\cdots&\tilde{Q}_{1N}\\ \vdots&\vdots&\cdots&\vdots\\ \tilde{Q}_{N1}&\tilde{Q}_{N2}&\cdots&\tilde{Q}_{NN}\end{bmatrix}$,
+with
+$\tilde{Q}_{ii}$ $=\sum_{j=1}^{N}c_{ij}(Q_{ij}+Q_{ji}),$
+$\tilde{Q}_{ij}$ $=-c_{ij}(Q_{ij}+Q_{ji}),$
+while $c_{ij}=1$ if $j\in N_{i}$, otherwise $c_{ij}=0$.
+In particular, each agent $i$ has access to the $i$-th block row of $\tilde{Q}$, i.e., $\tilde{Q}_{i}=L_{i}\tilde{Q}$ with $L_{i}=diag\{0,\cdots,0,NI,0,\cdots,0\}$.
+The objective of this section is to design a distributed controller by using its own and neighboring information of the system (48) to minimize the cost function (49).
+Remark 3
+In the literature, a commonly used protocol for multi-agent systems [37, 38, 39] is:
+$u_{i}(t)=K\sum_{j\in N_{i}}(x_{j}(t)-x_{i}(t)),$ (50)
+which ensures that the system (48) achieves state consensus, that is, for any initial state $x_{i}(0)$,
+$\lim_{t\rightarrow\infty}\|x_{i}(t)-x_{j}(t)\|=0,~~i,j=1,2,\cdots,N.$ (51)
+However, such protocols do not consider the optimality, and the optimality of some specific performance index, which is essential in practical scenarios such as smart grids [16] and unmanned ground vehicle (UGV) swarm coordination [32].
+In this section, we present an optimal consensus protocol using Algorithm 2.
+To address this, we first present the centralized optimal controller, which requires information from all agents.
+Lemma 6
+Assume that $(A_{i},B_{i})$ is stabilizable and $(A,\tilde{Q}^{1/2})$ is observable, where $A=diag\{A_{1},\cdots,A_{N}\}$. Then, the centralized optimal controller is given by
+$u^{\diamond}(t)$ $=-R^{-1}B^{\prime}Px^{\diamond}(t),$ (52)
+where $B=diag\{B_{1},\cdots,B_{N}\}$, $P$ is the unique positive-definite solution to the algebraic Riccati equation:
+$A^{\prime}P+PA+\tilde{Q}-PBR^{-1}B^{\prime}P$ $=0,$ (53)
+and $x^{\diamond}(t)$ satisfies:
+$\dot{x}^{\diamond}(t)$ $=[A-BR^{-1}B^{\prime}P]x^{\diamond}(t),\quad x^{\diamond}(0)=x(0).$ (54)
+Proof.
+Since $(A_{i},B_{i})$ is stabilizable, the block diagonal $(A,B)$ is stabilizable.
+Combined with the observability of $(A,\tilde{Q}^{1/2})$, the algebraic Riccati equation (53) admits a unique positive-definite solution $P$.
+The optimal controller (52) and the corresponding closed-loop system (54) then follow directly from standard LQR theory [34].
+This completes the proof.
+$\blacksquare$
+Next, we derive the distributed solution by using Algorithm 2. To this end, we define
+$\tilde{A}_{i}=diag\{0,\cdots,0,NA_{i},0,\cdots,0\},$
+$\tilde{B}_{i}=diag\{0,\cdots,0,NB_{i},0,\cdots,0\},$
+$\tilde{R}_{i}=diag\{0,\cdots,(1/N)R_{i}^{-1},\cdots,0\},$
+and introduce the distributed iteration equation for $P$:
+$0$ $=(\bar{Z}_{i,k}^{n})^{\prime}P_{i,k}^{n}+P_{i,k}^{n}\bar{Z}_{i,k}^{n}+\bar{V}_{i,k}^{n},$ (55)
+$\bar{Z}_{i,k}^{n}$ $=\bar{Z}_{i,k-1}^{n}+\alpha_{k}[\tilde{A}_{i}-\tilde{B}_{i}\tilde{R}_{i}\tilde{B}_{i}^{\prime}P_{i,k-1}^{n-1}-\bar{Z}_{i,k-1}^{n}]$
+$\quad+\frac{1}{\gamma}\sum_{j\in N_{i}}[\bar{Z}_{j,k-1}^{n}-\bar{Z}_{i,k-1}^{n}],$ (56)
+$\bar{V}_{i,k}^{n}$ $=\bar{V}_{i,k-1}^{n}+\alpha_{k}[\tilde{Q}_{i}+P_{i,k-1}^{n-1}\tilde{B}_{i}\tilde{R}_{i}\tilde{B}_{i}^{\prime}P_{i,k-1}^{n-1}$
+$\quad-\bar{V}_{i,k-1}^{n}]+\frac{1}{\gamma}\sum_{j\in N_{i}}[\bar{V}_{j,k-1}^{n}-\bar{V}_{i,k-1}^{n}],$ (57)
+with $\bar{Z}_{i,0}^{n}=0$, $P_{i,k}^{0}=0$, $P_{i,0}^{n}=0$, $\bar{V}_{i,0}^{n}=0$, and the distributed iteration equation for $x^{\diamond}(t)$:
+$x_{i,k}^{\diamond}(t)$ $=x_{i,k-1}^{\diamond}(t)+\alpha_{k}\big[e^{\bar{Z}_{i,\infty}t}\bar{x}_{i}(0)-x_{i,k-1}^{\diamond}(t)\big]$
+$\quad+\frac{1}{\gamma}\sum_{j\in\mathcal{N}_{i}}[x_{j,k-1}^{\diamond}(t)-x_{i,k-1}^{\diamond}(t)],$ (58)
+with $\bar{x}_{i}(0)=[0~\cdots~0~Nx_{i,0}^{\prime}~0~\cdots~0]^{\prime}$, $x_{i,0}^{\diamond}(t)=0$ and $\bar{Z}_{i,\infty}$ is the convergent value of (56).
+Then, we present the distributed design of the controller (52) as follows.
+Theorem 4
+Under the stabilizability of $(A_{i},B_{i})$ and observability of $(A,\tilde{Q}^{1/2})$, let $\alpha_{k}$ satisfy $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, and choose $\gamma>0$ such that $I_{N}-\frac{1}{\gamma}\mathcal{L}-\frac{1}{N}\mathbf{1}_{N}\mathbf{1}_{N}^{\prime}$ is Hurwitz, then it holds that:
+$P_{i,\infty}$ $\triangleq\lim_{n\to\infty}\lim_{k\to\infty}P_{i,k}^{n}=P,$~~
+$\bar{Z}_{i,\infty}$ $\triangleq\lim_{n\to\infty}\lim_{k\to\infty}\bar{Z}_{i,k}^{n}=A-BR^{-1}B^{\prime}P,$
+$x_{i,\infty}^{\diamond}(t)$ $\triangleq\lim_{k\to\infty}x_{i,k}^{\diamond}(t)=x^{\diamond}(t).$
+In this case, the distributed optimal controller for each agent $i$ is given by
+$u_{i}(t)=-[0~\ldots~0~R_{i}^{-1}B_{i}^{\prime}~0~\ldots~0]P_{i,\infty}x_{i,\infty}^{\diamond}(t),$ (59)
+which satisfies
+$\|u(t)-u^{\diamond}(t)\|\leq\varepsilon,$ (60)
+and ensures that the closed-loop system (48) achieves consensus.
+Proof.
+The proof is divided into two parts.
+First, we prove that (60) holds.
+Similar to the proof of Lemma 2, there exist integers $\bar{n}\geq n,\bar{k}\geq k$ and sufficiently small constants $\varepsilon_{1}>0$, $\varepsilon_{2}>0$ such that
+$\|P_{i,\bar{k}}^{\bar{n}}-P\|$ $\leq\varepsilon_{1},$ (61)
+$\|x_{i,\bar{k}}^{\diamond}(t)-x^{\diamond}(t)\|$ $\leq\varepsilon_{2}.$ (62)
+This implies that the distributed solutions $P_{i,k}^{n}$ and $x_{i,k}^{\diamond}(t)$ can approximate their global solution $P$ and $x^{\diamond}(t)$.
+Then, combining the fact that $u(t)=[u_{1}^{\prime}(t)~\cdots~u_{N}^{\prime}(t)]^{\prime}$ and
+$\begin{bmatrix}R_{1}^{-1}B_{1}^{\prime}&0&\cdots&0\\ 0&R_{2}^{-1}B_{2}^{\prime}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&R_{N}^{-1}B_{N}^{\prime}\end{bmatrix}=R^{-1}B^{\prime},$
+we can design the distributed controller (59) from (52) and obtain (60) holds.
+Second, we show that the system achieves consensus. Since the problem has a solution and $J<\infty$, then it follows from the properties of infinite integrals that
+$\lim_{t\rightarrow\infty}(x_{i}(t)-x_{j}(t))^{\prime}Q_{ij}(x_{i}(t)-x_{j}(t))=0.$ (63)
+Combined with $Q_{ij}>0$, the necessary condition for the term (63) to converge to zero over the infinite horizon is
+$\lim_{t\rightarrow\infty}\|x_{i}(t)-x_{j}(t)\|=0,~~i,j=1,2,\cdots,N.$
+This completes the proof. $\blacksquare$
+VI Simulation Example
 In this section, two examples are provided to show the effectiveness of the proposed Algorithms.
-Example 1. Consider the distributed optimal control of a heterogeneous UGVs system, where the dynamics of each UGV is given by [35]
+Example 1. Consider system (1) and the cost function (2) with parameters given by
+$A=\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right],B=\left[\begin{array}[]{cc}1\\ 1\end{array}\right],Q=I,R=1,$ (68)
+$x_{0}=\left[\begin{array}[]{cc}4\\ 4\end{array}\right],x_{T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right].$ (73)
+Moreover, 4 agents are involved in the sensor network and the topology graph is shown in Fig. 1.
+Fig. 1. Communication topology among agents.
+Next, we consider the distributed case where each agent only has access to partial system information. Based on Assumptions 1-4, the parameters for each agent are given by
+$A_{1}=\left[\begin{array}[]{cc}1.5&0\\ 0&0.5\end{array}\right],A_{2}=\left[\begin{array}[]{cc}0.5&0\\ 0&1.2\end{array}\right],$
+$A_{3}=\left[\begin{array}[]{cc}1.3&0\\ 0&1.8\end{array}\right],A_{4}=\left[\begin{array}[]{cc}0.7&0\\ 0&0.5\end{array}\right],$
+$B_{1}=\left[\begin{array}[]{cc}1.5\\ 0.5\end{array}\right],B_{2}=\left[\begin{array}[]{cc}0.5\\ 1.2\end{array}\right],B_{3}=\left[\begin{array}[]{cc}0.3\\ 1.8\end{array}\right],$
+$B_{4}=\left[\begin{array}[]{cc}1.7\\ 0.5\end{array}\right],Q_{1}=\left[\begin{array}[]{cc}1.2&0\\ 0&1\end{array}\right],Q_{2}=\left[\begin{array}[]{cc}1&0\\ 0&0.8\end{array}\right],$
+$Q_{3}=\left[\begin{array}[]{cc}1&0\\ 0&1.2\end{array}\right],Q_{4}=\left[\begin{array}[]{cc}0.8&0\\ 0&1\end{array}\right],R=1,$
+$x_{1,0}=\left[\begin{array}[]{cc}3.5\\ 4\end{array}\right],x_{2,0}=\left[\begin{array}[]{cc}4.5\\ 3.5\end{array}\right],x_{3,0}=\left[\begin{array}[]{cc}2.5\\ 4.5\end{array}\right],$
+$x_{4,0}=\left[\begin{array}[]{cc}5.5\\ 4\end{array}\right],x_{i,T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right],M_{i}=\left[\begin{array}[]{cc}1&0\\ 0&1\end{array}\right],$
+$i=1,\cdots,4.$
+By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, $\gamma=2.5$, $\varsigma=10^{-3}$, $n=20$, $k=\varpi=q=w=200$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figs.
+2-7.
+It is obvious that the distributed solutions converge to the centralized solutions from the Figs.
+2-7.
+Fig. 2. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
+Fig. 3. The trajectories of $\|W_{i,\varpi}\|_{2}$.
+Fig. 4. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
+Fig. 5. The trajectories of $\|\beta_{i}(t)\|_{2}$.
+Fig. 6. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$.
+Fig. 7. The trajectories of $u_{i}^{*}(t)$.
+Example 2.
+In the example, we will apply Theorem 4 to solve the consensus problem of a heterogeneous UGV system.
+Specifically, the objective is to enable five UGVs to achieve position consensus, where the dynamics of each UGV are given by [35]:
 $\left[\begin{array}[]{c}\dot{q}_{i}(t)\\ \dot{v}_{i}(t)\end{array}\right]=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right]\left[\begin{array}[]{c}q_{i}(t)\\ v_{i}(t)\end{array}\right]+\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right]F_{i}(t),$ (81)
-while $q_{i}(t)\in R^{2}$, $i=1,\cdots,5$, denote the position and velocity vectors of the $i$th UGV, $F_{i}(t)\in R^{2}$ denotes the force applied to the UGV, $D_{i}$ denotes the mechanical mass and $C_{i}$ denotes the translational friction coefficient.
-The communication topology among the UGVs is illustrated in Figure 2.
-Figure 2. Communication topology among heterogeneous UGVs.
-By defining the state variables as $x(t)=[x_{1}(t)^{\prime}~{}\cdots~{}x_{5}(t)^{\prime}]^{\prime}$ and the input variable as $u(t)=[u_{1}(t)^{\prime}~{}\cdots~{}u_{5}(t)^{\prime}]^{\prime}$, where ${x}_{i}(t)=[q_{i}(t)^{\prime}~{}~{}v_{i}(t)^{\prime}]^{\prime}$, ${u}_{i}(t)=F_{i}(t)$, then the UGV system (63) can be formulated as (3) with
-$A_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&5\tilde{A}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],B_{i}=\left[\begin{array}[]{cccc}0{}&{}\hfil&{}\hfil&\\ {}\hfil&5\tilde{B}_{i}{}&{}\hfil&\\ {}\hfil&{}\hfil&\ddots{}&\\ {}\hfil&{}\hfil&{}\hfil&0\\ \end{array}\right],$
-$\tilde{A}_{i}=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right],~{}~{}\tilde{B}_{i}=\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right].$
-The control objective is to drive the UGVs from initial positions to predefined target formation while minimizing the cost functions (2) with $Q=16I$, $R=10I$, the initial states of the UGVs being
-$x_{1}(0)=\left[\begin{array}[]{cccc}1.5\\ 2.5\\ 0\\ 0\end{array}\right],x_{2}(0)=\left[\begin{array}[]{cccc}2.5\\ 2.6\\ 0\\ 0\end{array}\right],x_{3}(0)=\left[\begin{array}[]{cccc}2\\ 1\\ 0\\ 0\end{array}\right],$
-$x_{4}(0)=\left[\begin{array}[]{cccc}0\\ 1\\ 0\\ 0\end{array}\right],x_{5}(0)=\left[\begin{array}[]{cccc}1\\ 2\\ 0\\ 0\end{array}\right].$
-and the terminal states of the UGVs being
-$x_{1}(T)=\left[\begin{array}[]{cccc}5.5\\ 8\\ 3\\ 4\end{array}\right],x_{2}(T)=\left[\begin{array}[]{cc}7\\ 7\\ 2\\ 4\end{array}\right],x_{3}(T)=\left[\begin{array}[]{cc}8\\ 6\\ 3\\ 6\end{array}\right],$
-$x_{4}(T)=\left[\begin{array}[]{cc}3\\ 6\\ 1\\ 3\end{array}\right],x_{5}(T)=\left[\begin{array}[]{cc}4\\ 7\\ 2\\ 1\end{array}\right].$
-Next, by running Algorithm 2 with step size $\alpha_{k}=\frac{1}{k}$ and $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 3-Figure 8.
-It is obvious that all the agents achieve consensus and converge to the centralized solutions from Figure 7-Figure 8.
-In particular, the trajectories of the agent achieve the predefined target formation at the given time $T=2$s as shown in Figure 9.
-Figure 3. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
-Figure 4. The trajectories of $\|W_{i,\varpi}\|_{2}$.
-Figure 5. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
-Figure 6. The trajectories of $\|\beta_{i}(t)\|_{2}$.
-Figure 7. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$ with $w=200$.
-Figure 8. The trajectories of $u_{i}^{*}(t)$.
-Figure 9. The predefined target formation of heterogeneous UGVs.
-Example 2. Consider system (1) and the cost function (2) with parameters given by
-$A=\left[\begin{array}[]{cc}1{}&2\\ 0{}&1\end{array}\right],B=\left[\begin{array}[]{cc}1\\ 1\end{array}\right],Q=I,R=1,$ (74)
-$x_{0}=\left[\begin{array}[]{cc}4\\ 4\end{array}\right],x_{T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right].$ (79)
-Moreover, 4 agents are involved in the sensor network and the topology graph is shown in Figure 10.
-Figure 10. Communication topology among agents.
-Next, we consider the distributed case where each agent only has access to partial system information. Based on Assumptions 1-4, the parameters for each agent are given by
-$A_{1}=\left[\begin{array}[]{cc}1.5{}&2\\ 0{}&0.5\end{array}\right],A_{2}=\left[\begin{array}[]{cc}0.5{}&2\\ 0{}&1.2\end{array}\right],A_{3}=\left[\begin{array}[]{cc}1.3{}&2\\ 0{}&1.8\end{array}\right],$
-$A_{4}=\left[\begin{array}[]{cc}0.7{}&2\\ 0{}&0.5\end{array}\right],B_{1}=\left[\begin{array}[]{cc}1.5\\ 0.5\end{array}\right],B_{2}=\left[\begin{array}[]{cc}0.5\\ 1.2\end{array}\right],$
-$B_{3}=\left[\begin{array}[]{cc}0.3\\ 1.8\end{array}\right],B_{4}=\left[\begin{array}[]{cc}1.7\\ 0.5\end{array}\right],Q_{1}=\left[\begin{array}[]{cc}1.2{}&0\\ 0{}&1\end{array}\right],$
-$Q_{2}=\left[\begin{array}[]{cc}1{}&0\\ 0{}&0.8\end{array}\right],Q_{3}=\left[\begin{array}[]{cc}1{}&0\\ 0{}&1.2\end{array}\right],Q_{4}=\left[\begin{array}[]{cc}0.8{}&0\\ 0{}&1\end{array}\right],$
-$R=1,x_{1,0}=\left[\begin{array}[]{cc}3.5\\ 4\end{array}\right],x_{2,0}=\left[\begin{array}[]{cc}4.5\\ 3.5\end{array}\right],x_{3,0}=\left[\begin{array}[]{cc}2.5\\ 4.5\end{array}\right],$
-$x_{4,0}=\left[\begin{array}[]{cc}5.5\\ 4\end{array}\right],x_{i,T}=\left[\begin{array}[]{cc}0\\ 0\end{array}\right],i=1,\cdots,4.$
-By running Algorithm 2 with $\alpha_{k}=\frac{1}{k}$, $\gamma=0.4$, we obtain $P_{i,k}^{n}(t)$, $W_{i,\varpi}$, $\lambda_{i,q}^{*}$, $\beta_{i}(t)$, $x_{i,w}^{*}(t)$ and $u_{i}^{*}(t)$, and the corresponding trajectories are shown in Figure 11-Figure 16.
-It is obvious that the distributed solutions converge to the centralized solutions Figure 15-Figure 16.
-Figure 11. The trajectories of $\|P_{i,k}^{n}(t)\|_{2}$.
-Figure 12. The trajectories of $\|W_{i,\varpi}\|_{2}$.
-Figure 13. The trajectories of $\|\lambda_{i,q}^{*}\|_{2}$.
-Figure 14. The trajectories of $\|\beta_{i}(t)\|_{2}$.
-Figure 15. The trajectories of $\|x_{i,w}^{*}(t)\|_{2}$.
-Figure 16. The trajectories of $u_{i}^{*}(t)$.
-6 Conclusions
+while $q_{i}(t)\in R^{2}$, $v_{i}(t)\in R^{2}$, $i=1,\cdots,5$, denote the position and velocity vectors of the $i$th UGV, $F_{i}(t)\in R^{2}$ denotes the force applied to the UGV, $D_{i}$ denotes the mechanical mass and $C_{i}$ denotes the translational friction coefficient.
+The system parameters are listed in Table I, and the communication topology among UGVs is illustrated in Fig.
+8.
+Fig. 8. Communication topology among heterogeneous UGVs.
+TABLE I: Parameters of the UGV system.
+Parameters UGV1 UGV2 UGV3 UGV4 UGV5
+$C_{i}$ $0.6$ $0.8$ $1.2$ 11 $0.4$
+$D_{i}$ 55 44 66 44 33
+By defining the state variables as $x_{i}(t)=[q_{i}(t)^{\prime}~~v_{i}(t)^{\prime}]^{\prime}$ and the control input as $u_{i}(t)=F_{i}(t)$, then the UGV system (81) can be formulated as (48) with
+$A_{i}=\left[\begin{array}[]{cc}0&I\\ 0&-\frac{C_{i}}{D_{i}}I\end{array}\right],~~B_{i}=\left[\begin{array}[]{c}0\\ \frac{1}{D_{i}}I\end{array}\right].$
+Let the parameters in cost function (49) be given by
+$Q_{11}=Q_{22}=Q_{33}=Q_{44}=Q_{55}=0,$
+$Q_{12}=Q_{15}=Q_{23}=Q_{34}=Q_{45}=I,$
+$Q_{21}=Q_{51}=Q_{32}=Q_{43}=Q_{54}=I,$
+$Q_{13}=Q_{14}=Q_{24}=Q_{25}=Q_{35}=0,$
+$Q_{31}=Q_{41}=Q_{42}=Q_{52}=Q_{53}=0,$
+$R_{1}=R_{2}=R_{3}=R_{4}=R_{5}=I.$
+In particular, the initial positions and velocities of the UGVs are given by
+$x_{1}(0)=\left[\begin{array}[]{cccc}2\\ 6\\ 1\\ 1\end{array}\right],x_{2}(0)=\left[\begin{array}[]{cccc}3\\ 3\\ 2\\ 2\end{array}\right],x_{3}(0)=\left[\begin{array}[]{cccc}2\\ 2\\ 1\\ 2\end{array}\right],$
+$x_{4}(0)=\left[\begin{array}[]{cccc}1\\ 2\\ 1.5\\ 1.5\end{array}\right],x_{5}(0)=\left[\begin{array}[]{cccc}1\\ 4\\ 2\\ 1\end{array}\right].$
+By selecting step size $\alpha_{k}=1/k$, $\gamma=2.5$, $\varepsilon=\varepsilon_{1}=\varepsilon_{2}=10^{-4}$, we execute the distributed iterative equations (55)-(58). Then it follows that:
+  • The distributed solutions $\|P_{i,\bar{k}}^{\bar{n}}\|_{2}$ with $\bar{n}=20$, $\bar{k}=900$ are:
+    $\|P_{1,\bar{k}}^{\bar{n}}\|_{2}=33.5490,~\|P_{2,\bar{k}}^{\bar{n}}\|_{2}=33.5134,$
+    $\|P_{3,\bar{k}}^{\bar{n}}\|_{2}=33.5116,~\|P_{4,\bar{k}}^{\bar{\bar{n}}}\|_{2}=33.5111,$
+    $\|P_{5,\bar{k}}^{n}\|_{2}=33.5475,$
+    which approximate the centralized solution $\|P\|_{2}=33.5591$ solved by (53).
+  • The trajectories of $x_{i,\bar{k}}^{\diamond}(t)$ for all UGVs are shown in Fig.
+9.
+It is obvious that $x_{i,k}^{\diamond}(t)$ converges to the centralized solution $x^{\diamond}(t)$ solved by (54).
+  • The trajectory of $u(t)$, composed of all distributed controllers $u_{i}(t)$, is shown in Fig.
+10, which demonstrates that $u(t)$ converges to the centralized optimal controller $u^{\diamond}(t)$ designed in (52).
+  • The position and velocity trajectories of the five UGVs are shown in Figs.
+11 and 12, respectively, which demonstrate that all UGVs achieve position consensus, and their velocities eventually converge to zero.
+In particular, we conduct a comparative simulation with the widely used classical consensus protocol [38] under the same initial conditions and system parameters $C_{i}$, $D_{i}$.
+The results are shown in Table II, from which it is obvious that the performance index $J$ of the proposed method is smaller than that obtained by using the controllers in [38].
+TABLE II: Performance index comparison between the proposed method and [38] under different $(C_{i},D_{i})$.
+Case $C_{i},D_{i}$ $J$ (Proposed Method) $J$ ([38])
+Case 1 (0.6, 5) 82.7573 125.1136
+Case 2 (0.8, 4) 76.2762 107.3754
+Case 3 (1.2, 6) 89.3999 135.6383
+Case 4 (1, 4) 76.4114 106.4391
+Case 5 (0.4, 5) 82.7004 127.5681
+Fig. 9. The trajectories of $\|x_{i,\bar{k}}^{\diamond}(t)\|_{2}$.
+Fig. 10. The trajectories of $\|u(t)\|_{2}$.
+Fig. 11. The trajectories of the position states $p_{i}(t)$ for all UGVs.
+Fig. 12. The trajectories of the velocity states $v_{i}(t)$ for all UGVs.
+VII Conclusions
 This paper studied the LQ optimal control problem for continuous-time systems with terminal state constraints.
 A distributed algorithm was proposed to solve the optimal controller where only partial information is accessible to the agent, consisting of the distributed iteration of the Riccati equation, the optimal Lagrange multiplier, and the optimal state.
-Numerical examples demonstrated the effectiveness of the proposed algorithm.
+As a significant application, the proposed distributed approach is extended to address the optimal consensus control problem.
+Numerical examples demonstrated the effectiveness of the proposed algorithm, where the performance index under the proposed distributed controller is smaller than that under the commonly used consensus control.
 Appendix A
+Proof of Lemma 2
 We will prove the result by using mathematical induction. For $n=1$, it is obvious from (21)-(22) that
 $Z_{i,k}^{1}(t)$ $=$ $Z_{i,k-1}^{1}(t)+\alpha_{k}[A_{i}-M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$ (85)
-$-Z_{i,k-1}^{1}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[Z_{j,k-1}^{1}(t)-Z_{i,k-1}^{1}(t)],$
+$-Z_{i,k-1}^{1}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[Z_{j,k}^{1}(t)-Z_{i,k-1}^{1}(t)],$
 $V_{i,k}^{1}(t)$ $=$ $V_{i,k-1}^{1}(t)+\alpha_{k}[Q_{i}+P_{i,k-1}^{0}(t)M_{i}M_{i}^{\prime}P_{i,k-1}^{0}(t)$ (86)
-$-V_{i,k-1}^{1}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[V_{j,k-1}^{1}(t)-V_{i,k-1}^{1}(t)].$
+$-V_{i,k-1}^{1}(t)]+\frac{1}{\gamma}\sum_{j\in N_{i}}[V_{j,k}^{1}(t)-V_{i,k-1}^{1}(t)].$
 It is aimed to show
-$\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z(t)\|$ $=$ $0,$ (88)
-$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V(t)\|$ $=$ $0$ (89)
+$\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z^{1}(t)\|$ $=$ $0,$ (87)
+$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0.$ (88)
 which is divided into two steps. The first step is to prove (85)-(86) achieves consensus, that is,
 $\displaystyle\lim_{k\rightarrow\infty}\|Z_{i,k}^{1}(t)-Z_{j,k}^{1}(t)\|$ $=$ $0,$ (89)
-$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V_{j,k}^{1}(t)\|$ $=$ $0$ (91)
+$\displaystyle\lim_{k\rightarrow\infty}\|V_{i,k}^{1}(t)-V_{j,k}^{1}(t)\|$ $=$ $0.$ (90)
 The second step is to prove (85)-(86) achieves global convergence, that is,
 $\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}Z_{i,k}^{1}(t)-Z^{1}(t)\|$ $=$ $0,$ (91)
-$\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0$ (93)
+$\displaystyle\lim_{k\rightarrow\infty}\|\frac{1}{N}\sum_{i=1}^{N}V_{i,k}^{1}(t)-V^{1}(t)\|$ $=$ $0.$ (92)
 Firstly, we prove that the consensus property by denoting
@@
 According to (85)-(86), we have that
-$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
+$\bar{\mathcal{Z}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{Z}}_{k-1}^{1}(t)$
 $-\alpha_{k}(U^{\prime}\otimes I_{n})\Pi_{k-1}^{1}(t),$
-$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
+$\bar{\mathcal{V}}_{k}^{1}(t)$ $=$ $[(I-\alpha_{k}I-\frac{1}{\gamma}U^{\prime}\mathcal{L}U)\otimes I_{n}]\bar{\mathcal{V}}_{k-1}^{1}(t)$
 $-\alpha_{k}(U^{\prime}\otimes I_{n})\Theta_{k-1}^{1}(t).$
@@
 $\delta_{k}^{1}(t)$ $=$ $[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\delta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}$
-×$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{\tau-1}^{1}(t),$
+$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Pi_{\tau-1}^{1}(t),$
 $\eta_{k}^{1}(t)$ $=$ $[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\eta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}$
-×$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{\tau-1}^{1}(t).$
+$\times[(I_{N}-\mathcal{B})\otimes I_{n}]\Theta_{\tau-1}^{1}(t).$
 This further gives that
@@
 $\leq$ $c\rho^{k}\|\delta_{0}^{1}(t)\|+\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
-×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|,$
+$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|,$
 $\|\eta_{k}^{1}(t)\|$ $=$ $\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k}\eta_{0}^{1}(t)+\sum_{\tau=1}^{k}\alpha_{\tau}[(\mathcal{A}-\mathcal{B})$
@@
 $\leq$ $c\rho^{k}\|\eta_{0}^{1}(t)\|+\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
-×$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|.$
+$\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|.$
 Since the graph is connected, there exists a constant $\gamma>0$ such that the following inequality holds
@@
 $\lim_{k\to\infty}\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
-×$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|=0,$ (98)
+$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Pi_{\tau-1}^{1}(t)\|=0,$ (97)
 $\lim_{k\to\infty}\sum_{\tau=1}^{k}\alpha_{\tau}\|[(\mathcal{A}-\mathcal{B})\otimes I_{n}]^{k-\tau}\|$
-×$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|=0.$ (99)
+$\quad\quad\quad\quad\quad\times\|[(I_{N}-\mathcal{B})\otimes I_{n}]\|\|\Theta_{\tau-1}^{1}(t)\|=0.$ (98)
 Thus, we obtain that
@@
 $\tilde{V}_{k}^{1}(t)$ $=$ $\tilde{V}_{k-1}^{1}(t)+\frac{1}{N}\alpha_{k}\sum_{i=1}^{N}[Q_{i}+P_{i,k-1}^{0}(t)M_{i}M_{i}^{\prime}$
-×$\times P_{i,k-1}^{0}(t)-V_{i,k-1}^{1}(t)].$
+$\times P_{i,k-1}^{0}(t)-V_{i,k-1}^{1}(t)].$
 By defining $Z_{k}^{1}(t)$, $V_{k}^{1}(t)$ by
@@
 $V_{k}^{n}(t)$ $=$ $V_{k-1}^{n}(t)+\alpha_{k}[Q+P^{n-1}(t)BR^{-1}B^{\prime}$ (100)
-×$\times P^{n-1}(t)-V_{k-1}^{n}(t)],$
+$\times P^{n-1}(t)-V_{k-1}^{n}(t)],$
 and letting $\Delta Z_{k}^{1}(t)=\tilde{Z}_{k}^{1}(t)-Z_{k}^{1}(t)$, $\Delta V_{k}^{1}(t)=\tilde{V}_{k}^{1}(t)-V_{k}^{1}(t)$, we have that
@@
 $\Delta{V}_{k}^{1}(t)$ $=$ $(1-\alpha_{k})\Delta{V}_{k-1}^{1}(t).$
-By applying the facts that $10<1-\alpha_{k}<$ 1, and $\lim_{k\to\infty}\alpha_{k}=0$, it yields that
+By applying the facts that $0<1-\alpha_{k}<1$, and $\lim_{k\to\infty}\alpha_{k}=0$, it yields that
 $\lim_{k\to\infty}\|\Delta{Z}_{k}^{1}(t)\|$ $=$ $0,$ (101)
-$\lim_{k\to\infty}\|\Delta{V}_{k}^{1}(t)\|$ $=$ $0$ (103)
+$\lim_{k\to\infty}\|\Delta{V}_{k}^{1}(t)\|$ $=$ $0.$ (102)
 Combining with (89)-(90), it follows that
 $\lim_{k\to\infty}\|Z_{i,k}^{1}(t)-Z_{k}^{1}(t)\|$ $=$ $0,$
-$\lim_{k\to\infty}\|V_{i,k}^{1}(t)-V_{k}^{1}(t)\|$ $=$ $0$
+$\lim_{k\to\infty}\|V_{i,k}^{1}(t)-V_{k}^{1}(t)\|$ $=$ $0.$
 Moreover, together with the fact that $\lim_{k\rightarrow\infty}\|Z_{k}^{1}(t)-Z^{1}(t)\|=0$ and $\lim_{k\rightarrow\infty}\|V_{k}^{1}(t)-V^{1}(t)\|=0$ obtained by using Theorem 3.1.1 in [36], we obtain (87)-(88) directly.